### v1

# ChatDBG: An AI-Powered Debugging Assistant

Abstract
This paper presents ChatDBG, the first AI-powered debugging assistant. ChatDBG integrates large language models (LLMs) to significantly enhance the capabilities and user-friendliness of conventional debuggers. ChatDBG lets programmers engage in a collaborative dialogue with the debugger, allowing them to pose complex questions about program state, perform root cause analysis for crashes or assertion failures, and explore open-ended queries like why is x null?. To handle these queries, ChatDBG grants the LLM autonomy to take the wheel and drive debugging by issuing commands to navigate through stacks and inspect program state; it then reports its findings and yields back control to the programmer. Our ChatDBG prototype integrates with standard debuggers including LLDB, GDB, and WinDBG for native code and Pdb for Python. Our evaluation across a diverse set of code, including C/C++ code with known bugs and a suite of Python code including standalone scripts and Jupyter notebooks, demonstrates that ChatDBG can successfully analyze root causes, explain bugs, and generate accurate fixes for a wide range of real-world errors. For the Python programs, a single query led to an actionable bug fix 67% of the time; one additional follow-up query increased the success rate to 85%. ChatDBG has seen rapid uptake; it has already been downloaded nearly 30,000 times.
11footnotetext: Equal contribution.footnotetext: Work done at the University of Massachusetts Amherst.
I Introduction
Debuggers help programmers identify and fix bugs by letting them investigate program state and navigate program execution. Debuggers for mainstream languages, including GDB [1], LLDB [2], and WinDBG (for C, C++, and Rust), jdb (Java), Pdb (Python), and the Chrome or Firefox debuggers (for JavaScript), generally provide the same functionality. In particular, most debuggers support observing program execution via tracing and reporting when a program reaches a given line or function of source code; interrupting execution and returning control to the debugger when the program reaches a given line or function via breakpoints, when a particular condition is true via conditional breakpoints, or when a variable changes via watchpoints (a.k.a. data breakpoints); inspecting local variables, globals, heap objects, and backtraces of the call stack; resuming program execution line-by-line (single-step) or at the granularity of function calls; and in some debuggers, stepping backward through execution via reverse debugging, also known as time-travel or omniscient debugging.
Debuggers can be helpful, but finding and fixing software defects remains a deeply challenging and time-consuming task. Programmers must still reason about program behavior to ascertain what went wrong. They must formulate and test hypotheses about program execution, they must read and understand code they may have not written, and they must pore over potentially voluminous information. Such information includes lengthy executions, large amounts of program data, and many stack frames that potentially span multiple threads.
This paper introduces ChatDBG, the first AI-powered debugger assistant. ChatDBG integrates into and significantly extends the functionality of standard debuggers. ChatDBG builds on the insight that large language models (LLMs), such as OpenAI’s GPT-4 [3], enable a debugger to leverage insights and intuition from many thousands of programs as well as the vast real-world knowledge embedded in LLMs.
A debugger integrated with ChatDBG continues to provide its full range of functionality, but also lets programmers engage in debugging dialogs where they can ask high-level questions like why is x null here? or why isn’t this value what I expected?. The question can be as simple as why? if a program has crashed or failed an assertion. ChatDBG then orchestrates a conversation with the LLM. A key novelty of ChatDBG is that it grants autonomy to the LLM to “take the wheel” while answering the programmer’s queries. Specifically, the LLM issues “function calls” [4] to run commands in the underlying debugger to investigate program state, execute code, or obtain source code. The results of those calls are sent back to the LLM for use while constructing its response. After answering a query, control is returned to the programmer, who may then enter additional commands or chat messages.
Our prototype of ChatDBG integrates into four widely used debuggers: GDB, LLDB, WinDBG, and Pdb. Our evaluation presents a range of case studies demonstrating that ChatDBG improves significantly on existing debuggers. On a suite of unpublished Python scripts and Jupyter notebooks written by undergraduate students, one or two queries is sufficient for ChatDBG to properly diagnose and fix defects 87% of the time, typically at a cost of under \$0.20 USD. ChatDBG is also effective at identifying causes and providing fixes for a range of real-world bugs in C/C++ code.
This paper makes the following contributions: it introduces ChatDBG, the first AI-powered debugger assistant; it describes the implementation of our ChatDBG prototype; and it introduces the “take the wheel” approach to integrating large language models with developer tools. The paper also presents an evaluation of ChatDBG that demonstrates its significant advantages over existing debugger functionality.
This paper is organized as follows. Section II illustrates the use of ChatDBG to debug a program. Section III describes key related work. Section IV describes ChatDBG’s implementation. Section V presents our evaluation, and Section VI concludes with several promising directions for future work.
Source code for bootstrap.py
from datascience import *
from ds101 import *
def make_marble_sample():
    table = Table().read_table(’marble-sample.csv’)
    return table.column(’color’)
def proportion_blue(sample):
    return sample
def resampled_stats(observed_marbles, num_trials):
    stats = bootstrap_statistic(observed_marbles,
                                proportion_blue,
                                num_trials)
    assert len(stats) == num_trials
    return stats
observed_marbles = make_marble_sample()
stats = resampled_stats(observed_marbles, 5)
assert np.isclose(np.mean(stats), 0.7)
Figure 1: An example program containing several bugs. It is supposed to create an array of marble colors, compute the proportions of blue marbles in resamples of that array, and assert that their mean is about 0.7, the proportion for the array.
\$ chatdbg bootstrap.py Traceback (most recent call last):   File "/python3.11/site-packages/ipdb/__main__.py", ...    pdb._run(stdlib_pdb._ScriptTarget(mainpyfile))  File "/python3.11/pdb.py", line 1659, in _run    self.run(target.code)  File "/python3.11/bdb.py", line 600, in run    exec(cmd, globals, locals)  File "<string>", line 1, in <module>  File "./bootstrap.py", line 19, in <module>    stats = resampled_stats(observed_marbles, 5)            ^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^  File "./bootstrap.py", line 15, in resampled_stats    assert len(stats) == num_trials           ^^^^^^^^^^^^^^^^^^^^^^^^AssertionErrorUncaught exception. Entering post mortem debuggingRunning ’cont’ or ’step’ will restart the program (ChatDBG) p num_trials 5 (ChatDBG) Why doesn’t stats have 5 elements? (ChatDBG) p len(stats) 150 (ChatDBG) info bootstrap_statistic bootstrap_statistic(observed_sample,                     compute_statistic, num_trials)    Creates num_trials resamples of the initial    sample. Returns an array of the provided    statistic for those samples.    * observed_sample: the initial sample, as an                       array.    * compute_statistic: a function that takes a                         sample as an array and                         returns the statistic for                         that sample.    * num_trials: the number of bootstrap samples                  to create.You MUST assume that ‘bootstrap_statistic‘ isspecified and implemented correctly. (ChatDBG) p proportion_blue <function proportion_blue at 0x1050b00e0> (ChatDBG) info proportion_blue   8     def proportion_blue(sample):   9         return sample From the information we have, the stats variable in resampled_stats is supposed to have 5 elements, but it actually has 150 elements… The Python debugger output shows a reference to a variable or function called proportion_blue, which is used as the compute_statistic parameter for bootstrap_statistic. Upon looking at the proportion_blue function’s definition, it is found that the function simply returns the input sample without computing a statistical value. This means that for each of the num_trials (which is 5), bootstrap_statistic is adding the entire sample to the stats array without processing the sample to obtain a single statistic value…
Recommendation
To fix this issue, the proportion_blue function must be corrected to perform the intended computation of returning the proportion of blue marbles in the sample: def proportion_blue(sample):    return np.mean(sample == 'B')
Figure 2: A debugging session with ChatDBG. For brevity, we elide repetitive or unimportant parts. ChatDBG catches the assertion failure on line 15 and prompts the user to enter commands (bold and highlighted in gray). Commands issued by ChatDBG when it takes the wheel are highlighted in yellow, and its response is highlighted in blue.
II Overview
This section illustrates ChatDBG’s features and ability to assist in debugging the program in Figure 1. That program is a distillation of real errors encountered by students in an introductory data science lab. It creates an array observed_marbles representing the colors of marbles (red or blue) in a sample stored in a file. It then calls bootstrap_statistic to create same-sized resamples of that array. That function computes a statistic for each resample and returns an array of those statistics. In this case, the statistic is proportion_blue, the proportion of blue marbles. Given a sufficiently large number of trials, the mean of the resamples’ statistics should be close to 0.7, the proportion of blue marbles in the original sample [5].
The program fails the assertion in resampled_stats, and Figure 2 illustrates a debugging session. To try to figure out what went wrong, the user issues the Pdb command p num_trials to view the value of that variable. Continuing debugging with existing tools would likely involve issuing additional commands, examining data files, source code, and examining library documentation. With ChatDBG, the user instead starts a dialog with the debugger, asking why doesn’t stats have 5 elements? While constructing the answer (in blue), the LLM takes the wheel and directly issues debugger commands (yellow). These include standard Pdb commands and a ChatDBG-specific info command for accessing the source code and docstrings for any user-written code, and the docstrings for library code (which we assume is correct and not the root cause of any error).
\$ chatdbg bootstrap.py Traceback (most recent call last):   File "./bootstrap.py", line 21, in <module>    assert np.isclose(np.mean(stats), 0.7)           ^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^AssertionError... (ChatDBG) Why is the mean of stats not close to 0.7? (ChatDBG) p np.mean(stats) 0.7466666666666667 (ChatDBG) info resampled_stats ... (ChatDBG) info make_marble_sample .... (ChatDBG) info bootstrap_statistic ...
Recommendation
1. Check the CSV file marble-sample.csv to verify the proportion of blue to red marbles and ensure it is close to 70% blue marbles. 2. Review the implementation of the proportion_blue function… 3. Verify that the proportion_blue function … is available in the current scope or imported correctly.
(ChatDBG) p proportion_blue(observed_marbles) 0.7 (ChatDBG) That value is correct. (ChatDBG) p stats array([0.66666667, 0.8       , 0.63333333,        0.76666667, 0.86666667]) At this point, the most straightforward explanation for the error is that the number of trials is not enough to accurately reflect the expected mean due to random sampling effects. Increasing the number of trials would likely result in an average statistic that is closer to the expected value of 0.7, following the Law of Large Numbers.
Recommendation
Increase the number of bootstrap trials significantly (e.g., from 5 to something like 1000 or more) to reduce the variability of the mean…
Figure 3: A debugging session demonstrating how ChatDBG incorporates real-world knowledge. After correcting proportion_blue in Figure 1, the program fails on line 21 because the mean proportion of blue marbles in the resamples is not the expected value. ChatDBG identifies high variance resulting from the small number of trials as the root cause.
ChatDBG identifies the root cause: proportion_blue fails to compute and return the desired statistic, and it provides a corrected version of proportion_blue. When ChatDBG cannot identify the root cause, it suggests further debugging steps and control is returned to the user, who may continue the chat, issue further debugger commands, or both. Figure 3 illustrates this scenario, where a version of bootstrap.py with the corrected proportion_blue function fails the assertion on line 21. The user asks why the mean of stats is not close to 0.7, and ChatDBG’s first response suggests examining whether 0.7 is the appropriate expected value. The user then computes the proportion of blue marbles with a debugger command and tells ChatDBG that 0.7 is the correct value. In response, ChatDBG points to the low number of trials (5) as the issue. The LLM drew this correct conclusion without seeing any discussion of trial size or variance in any program state, code, or documentation encountered during the chat. A powerful aspect of ChatDBG is its ability to exploit real-world knowledge in its analyses (here, the fact that bootstrapping depends on large numbers of resamples) without specific instruction or user intervention.
III Related Work
System & Date Single Stack Break- Cond. Source Trace Display Eval. Watch- Reverse Explain Propose Open
Introduced Step Nav. points " Level State Code points Bugs Fixes Queries
DDT [6], 1961
EXDAMS [7], 1969
Mesa [8], 1979 ✓*
DBX [9], 1981 1990
GDB [10], 1986 1991 2009
Pdb, 1992
WinDBG, ca. 1997 2017
LLDB, 2010
ChatDBG, 2023
TABLE I: Debugger features and their dates of introduction. Most key features have been around for decades. By integrating into modern debuggers (GDB, LLDB, Pdb and WinDBG), ChatDBG inherits all of their features while significantly extending them with functionality to explain bugs and their root causes, propose fixes, and answer arbitrary natural-language queries over program state. (An asterisk or year in italics means the feature is limited in functionality, performance, or depends on specific hardware support; for example, WinDBG’s reverse debugging only works on specific Intel chips.)
Table I presents an overview of previous interactive debuggers, together with their features and date of introduction. The first interactive debugger, DDT, introduced breakpoints, single-stepping, and stack navigation in 1961 [6]. By 1979, the Mesa debugger had most key features of modern debuggers, including source-level debugging, conditional breakpoints, tracing, and the ability to display runtime state and evaluate code [8]. Arbitrary conditional breakpoints date back at least to 1990 with DBX [9]. Watchpoints were introduced by 1991 and have been in GDB since version 4.0.1 [10]. Reverse debugging it was first implemented in EXDAMS in 1969 [7] is present in widely-used debuggers like GDB (in 2009) [11], and WinDBG (in 2017); the rr debugger, built on top of GDB, also supports reverse debugging on Linux platforms [12]. ChatDBG, by integrating into standard debuggers, inherits and extends their functionality.
Ko and Myers present Whyline, an interactive, trace-based debugger that lets programmers select from a range of queries and identifies (via static and dynamic analysis) a timeline that answers the query [13]. Programmers can only select from those queries presented by Whyline as options. In contrast, ChatDBG permits programmers to pose arbitrary queries that it answers via a dialog with an LLM. Whyline’s use of traces gives it the ability to answer questions that might not be straightforward to answer with the current program state but limits its applicability to relatively short-lived executions.
The goal of program slicing, introduced by Weiser in 1981 [14], is to produce a shorter version of a program limited to the source code that could have led to an error. Program slicing has been extensively studied; Weiser’s paper has been cited over 5,000 times to date. As Section IV-G describes, ChatDBG performs backwards slicing to collect code spread across code cells to facilitate debugging of Jupyter notebooks.
Fault localization seeks to identify the likely location of the root cause of a defect; Wong et al. present a survey with over 400 citations [15]. Unlike previous work, ChatDBG performs fault localization by leveraging LLM-based examination and LLM-driven exploration of source code and program state.
Automated program repair is another highly active area of software engineering research [16]; its goal is to generate source-level program patches that prevent a program from failing. Unlike past work, ChatDBG performs best-effort automated program repair—leveraging its integration within a debugger—via LLM-based examination and LLM-driven exploration of program source and state.
IV Implementation
[…]
1:history = “”
2:loop
3:    line = Input()
4:    if IsDebuggerCommand(line) then
[…]
6:        Print(output)
7:        history = history + (line + $\rightarrow$” + output)
8:    else
[…]
10:             prompt = MakePrompt(Instructions(),
11:                                    EnrichedStack(), Inputs(),
12:                                    Error(), history, line)
13:         else
14:             prompt = MakePrompt(history, line)
[…]
16:         history = “”
17:         while ResponsePending() do
18:             match Receive()
[…]
20:                      output = DoCommand(cmd)
21:                      Print (cmd + $\rightarrow$” + output)
22:                      Send(output)
[…]
Figure 4: ChatDBG architecture and top-level command processing loop.
IV-A Using ChatDBG: Preliminaries
ChatDBG integrates with existing debuggers as either a plug-in or a direct extension. Our primary focus to date has been an extension to Pdb, which supports both non-interactive Python scripts and interactive sessions in IPython or Jupyter notebooks, and a plug-in for LLDB to support C/C++ code. A subset of features has been ported to GDB and WinDBG.
Configuration for Python is minimal and limited to the installation of the chatdbg package with the standard package installer, plus one optional shell script command to add it as an extension to IPython. ChatDBG extends either the standard pdb.Pdb debugger or IPython’s implementation of Pdb, depending on how it is run. Configuration for LLDB and other C/C++ debuggers is similarly straightforward. LLDB can be installed through standard package managers if it is not already present, and the ChatDBG plug-in is installed via a single shell command. Since ChatDBG leverages OpenAI’s LLMs, the user must also set an environment variable to a valid OpenAI API key within their system’s configuration settings.
Instructions:
  You are a debugging assistant. You will be given a Python stack trace for an error and answer questions related to the root cause of the error. Call the debug function to run Pdb debugger commands on the stopped program. You may call the debug function to run the following commands: bt, up, down, p expression, list. Call debug to print any variable value or expression that you believe may contribute to the error. Call the info function to get the documentation and source code for any variable, function, package, class, method reference, field reference, or dotted reference visible in the current frame. Examples include: n, e.n where e is an expression, and t.n where t is a type. Unless it is from a common, widely-used library, you MUST call info exactly once on any symbol that is referenced in code leading up to the error. Call the provided functions as many times as you would like. The root cause of any error is likely due to a problem in the source code from the user. Explain why each variable contributing to the error has been set to the value that it has. Continue with your explanations until you reach the root cause of the error. Your answer may be as long as necessary. End your answer with a section titled “Recommendation” that contains one of: a fix if you have identified the root cause a numbered list of 1-3 suggestions for how to continue debugging if you have not
Enriched Stack Trace:
  The program has this stack trace: [... skipping 4 hidden frame(s)] ./bootstrap.py(19)<module>()      15      16     assert len(stats) == num_trials      17      18 observed_marbles = make_marble_bag() ---> 19 resampled_stats(observed_marbles, 5)    Global variables:      observed_marbles: ndarray = array([’R’, R’, R’,        ..., B’, B’, B’], dtype=’<U1’) > ./bootstrap.py(16)resampled_stats()      14                                 num_trials)      15 ---> 16     assert len(stats) == num_trials      17      18 observed_marbles = make_marble_bag()    Variables in this frame:      num_trials: int = 5      observed_marbles: ndarray = array([’R’, R’, R’,        ..., B’, B’, B’], dtype=’<U1’)      stats: ndarray = array([’B’, R’, B’, ..., R’,        B’, R’], dtype=’<U32’)
Error:
  The program encountered the following error: AssertionError The code assert len(stats) == num_trials is correct and must not be changed.
History:
  This is the history of some pdb commands I ran and the results:   (ChatDBG) p num_trials   5
User Text:
  Why doesn’t stats have 5 elements?
Figure 5: The initial prompt for the debugging session in Figure 2. For brevity, the enriched stack includes only five lines of source in each frame, rather than the default of 10.
IV-B Debugging a Target Program
For Python, debugging with ChatDBG begins by running chatdbg on the target program. No special preparation of the target is needed; Python’s managed run time ensures that debugging information and source code is always available. Debugging is supported in IPython interactive sessions or Jupyter notebooks via the standard command-line flag --pdb or the Jupyter magic command %pdb, respectively. Control drops into the debugger when an exception occurs.
For C and C++, debugging begins by running lldb on the target program. The target program must be an unstripped executable generated with the -g compiler flag, which ensures the availability of DWARF debug information that describes the memory layout and maps the program’s machine code back to the original source code. That information is essential for the effective debugging of unmanaged code.
ChatDBG also handles native code generated for other languages but may require additional steps. For example, to debug a Rust target program, the Cargo.toml file must list ChatDBG as a dependency and the main function must be annotated with #[chatdbg::main] to ensure that error messages are visible to ChatDBG through a log file.
IV-C ChatDBG Architecture Overview
ChatDBG orchestrates communication between the user, the debugger, as shown in the architecture diagram and ChatDBG’s top-level command processing loop in Figure 4. The operations in the pseudocode map naturally onto debugger APIs and onto LLM APIs supporting completion and function calls [4]. ChatDBG currently utilizes OpenAI’s API [17] and GPT-4 models. We elaborate on the most salient technical innovations after this overview.
[…]
Command Debugger Output
info symbol Pdb The source code and/or docstring for a symbol referring to any function, method, field, class, or package.
code loc LLDB The source code surrounding loc, where loc is filename:lineno.
definition loc symbol LLDB The declaration for the first occurrence of symbol at loc, where loc is filename:lineno.
slice symbol Pdb The source code in the backwards slice of the global symbol. Interactive IPython/Notebook sessions only.
TABLE II: ChatDBG command extensions. These commands are available to not only the user but also the LLM, and they provide access to information beyond what is typical in a debugger.
IV-D Initial Prompts and Enriched Stack Traces
In addition to including the user’s text, the initial prompt conveys instructions to LLM and also the context surrounding the error. We highlight the components comprising that context in this section, using the initial prompt in Figure 5 that was generated for the first query in Figure 2.
Instructions. The instructions at the top of the prompt ask the LLM to answer questions about the root cause of the error, to focus on user code, to explain values stored in variables, and to end each response with either a fix or suggestions for further debugging steps. The last item ensures a relatively consistent structure on answers that facilitates reading them and evaluating their quality. Paragraphs 2-4 of the instructions are the take the wheel prompt described in Section IV-E.
Enriched stack trace. ChatDBG’s success at identifying and fixing errors relies critically on providing the LLM with sufficient details to reveal the cause of the error. A key source of that information is the run-time stack. Debuggers provide a way for the user to view the stack trace but often only show function names, source file locations, and possibly a couple lines of code for each stack frame. ChatDBG provides a more detailed enriched stack trace to the LLM. That stack trace includes the types and values of variables for each frame, as well as a larger window of at least 10 lines of code. Enriched stack traces also elide frames corresponding to library code to better focus the LLM on user-written code, which ChatDBG assumes to be the most likely cause of errors.
In Python, ChatDBG leverages Pdb’s internal data structures to build enriched stack traces. When converting values to suitable string representations, ChatDBG must balance utility with the size of the string produced. For objects, ChatDBG calls the object’s __repr__ method if an appropriate (non-default) version exists. Otherwise, it iterates over the object’s fields and recursively converts their values to strings. Similarly, ChatDBG recursively converts the values stored in aggregate structures like lists, arrays, and dictionaries to strings, but limits the number of elements shown to a small, fixed number. The rest of the elements are just abbreviated with an ellipsis (…). This recursive conversion of values to strings is limited to a depth of three, at which point any remaining values are again abbreviated with ellipses.
ChatDBG follows roughly the same approach in LLDB, utilizing the static types embedded in the DWARF debugging information to decode the stack. In addition, any pointers are dereferenced to show the values being referred to as well; null pointers and illegal dereferences are dropped.
Inputs. The initial prompt also includes the target’s command line arguments and standard input when that information is available from the underlying debugger. These are empty and elided in Figure 5.
Error. A description of the error causing execution to stop is also extracted from the underlying debugger. When the error is due to an assertion failure, ChatDBG instructs the LLM to assume that the assertion is valid as written so that it will look beyond the assertion for the real problem.
History. The initial prompt also includes the history of commands already issued by the user, as well as their outputs. This builds a more complete context surrounding the user’s query.
Name LoC Type Reported Exception Root Cause
c1 48 semantic Assertion Error Off-by-one error in an h-index computation
c2 81 crash Name Error Parameter not referenced properly
c3 64 crash Value Error Error in CSV column label leads to improper data parsing
c4 89 crash Index Error A class’s __str__ fails if an object’s internal list is empty
c5 29 crash Index Error Missing one of two base cases in a recursive function
c6 72 crash Name Error Multiple errors related to building list of user-defined objects
c7 71 semantic Assertion Error Failure to convert input to lower case before processing
c8 72 semantic Assertion Error Missing test for lowercase words
s1 123 semantic Assertion Error Incorrect drop and rename operations leading to bad data in table
s2 124 semantic Assertion Error Incorrect max operation on a table
s3 124 semantic Assertion Error Incorrect aggregation function in pivot operation
s4 124 semantic Assertion Error Incorrect aggregation function in group operation
s5 162 semantic Assertion Error Hardcoded table data in wrong order
s6 162 crash Name Error Typo in variable reference
s7 45 semantic Assertion Error Function confuses parameter and global variable
s8 49 semantic Assertion Error Wrong percentile used in confidence interval construction
s9 112 semantic Assertion Error Wrong percentile used in confidence interval construction
s10 118 semantic Assertion Error Loops doesn’t append to array correctly
s11 181 crash Value Error Takes a sample from a table smaller than the sample size without replacement
s12 127 crash Value Error Incorrect label when accessing column value for table row
s13 127 crash Value Error Pivot uses wrong columns for row/columns in new table
s14 65 crash Index Error Simulation fails to properly create a random sample under null hypothesis
TABLE III: Python programs exhibiting a variety of common errors. Programs c1–c8 are command line scripts, and programs s1–s14 are Jupyter notebooks, which utilize two non-standard libraries consisting of 3,000 lines of code. Semantic errors reflect failed tests expressed as assertions. Crashes reflect unexpected termination due to any other type of error.
IV-E Taking the Wheel
ChatDBG supports take the wheel debugging via the function call capabilities in OpenAI’s API and most recent models [4]. This feature lets clients register callback functions with the LLM for obtaining additional information while constructing a response. The LLM calls these functions by sending special messages to the client as part of its response stream. The client receives those messages, computes the requested results, and sends them back to the LLM. The initial prompt describes how to use the available functions.
For example, ChatDBG registers a debug(command) function for running a command in the underlying debugger. The LLM calls debug("p len(stats)") through this mechanism in the session from Figure 2. ChatDBG then runs Pdb’s command processing routine, onecmd("p len(stats)"), and captures the output to and send back. ChatDBG similarly uses the SBCommandInterpreter.HandleCommand routine in LLVM. In both cases, the command and output are printed so the user can see these steps.
The LLM has sufficient background knowledge on debuggers and requires no additional training to navigate up/down the stack, inspect variables and heap data, evaluate expressions, and perform other typical debugger operations.
IV-F Navigating the Code
While the LLM can often leverage pre-existing background knowledge of common Python and C/C++ standard libraries, it will likely have limited-to-no knowledge of any user-defined code or third-party library functions. ChatDBG extends the underlying debuggers with several new commands that are designed to help the LLM navigate through and understand the target’s code. These commands are available to the LLM via function calls and are listed in Figure II.
ChatDBG augments Pdb with the info command, which prints the docstring for any function, class, field, method, or package. It additionally prints the source for any user-defined code. The info requests in Figure 2 demonstrate these two cases for proportion_blue and bootstrap_statistic, respectively. The command is implemented via the standard inspect and pydoc libraries.
The info command is not directly reproducible for unmanaged code in LLVM because there is no comparable existing debugger support for retrieving the source or documentation for a symbol. Instead, ChatDBG adds two other debugging commands to LLDB. The first, code, prints the code surrounding a source location described by a filename and line number, as in code polymorph.c:118. The second command, definition, prints the location and source code for the definition corresponding to the first occurrence of a symbol on a given line of code. For example, definition polymorph.c:118 target prints the location and source for the declaration of target corresponding to its use on that line. The definition implementation leverages the clangd language server, which supports source code queries via JSON-RPC and Microsoft’s Language Server Protocol [18].
IV-G Slices for Interactive Python
ChatDBG supports debugging interactive IPython sessions and Jupyter notebooks. Interactive sessions lead to many individual code cells that are each evaluated separately. Cells may be evaluated out-of-order, override definitions from earlier cells, and communicate values to other cells through top-level global variables. Others have noted the challenges of reasoning about program behavior in this context [19, 20]. ChatDBG provides an additional slice debugging command to facilitate that reasoning. The slice command computes the backwards slice for any variable used in the current cell that was defined in previously-executed cells. It returns the code for cells in that slice. Suppose the code from bootstrap.py in Figure 1 were written in four notebook cells:
In[2]: def make_marble_sample(): ...
In[3]: def proportion_blue(sample): ...
In[4]: def resampled_stats(observed_marbles, num_trials):
           stats = bootstrap_statistic(observed_marbles,
                                       proportion_blue,
                                       num_trials)
           assert len(stats) == num_trials
           return stats
In[5]: observed_marbles = make_marble_sample()
       stats = resampled_stats(observed_marbles, 5)
After evaluating these cells, slice(observed_samples) returns the source for the cells labeled In[2] and In[5], and slice(stats) returns the source for all four cells. ChatDBG uses ipyflow to compute slices [20, 21].
V Evaluation
We demonstrate ChatDBG’s capacity to identify the root cause of defects and provide fixes in two contexts: bugs in relatively small Python programs written by students and bugs in large C/C++ programs. The former have well-defined expected behavior that enable us to thoroughly and systematically assess ChatDBG. The latter demonstrates its effectiveness on unmanaged code when unusual corner cases trigger crashes.
Our evaluation addresses the following research questions: RQ1: Is ChatDBG effective at diagnosing and fixing bugs in Python? RQ2: Which components of ChatDBG contribute to its effectiveness? RQ3: Is ChatDBG effective at diagnosing and fixing bugs in unmanaged code (C/C++)?
V-A Python
We applied ChatDBG to bugs in a collection of student labs from two introductory computer science courses; see Table III. Bugs c1–c8 are in non-interactive scripts from a programming class that perform various file reading and text processing tasks. Bugs s1–s14 are in Jupyter notebooks [22] from a data science class that manipulate, visualize, and compute over arrays and tables. Some bugs were apparent to the programs’ authors. Others were identified during autograding.
Unlike many existing bug benchmarks for Python, these programs are unpublished and thus not in the language model’s training data. In addition, the programs have clear correctness criteria that lead to objective effectiveness metrics in our experiments. The bugs are also representative of mistakes often made in languages like Python. They range from scoping issues, algorithmic errors, and misuse of library functions to subtle misunderstanding of domain knowledge. They include both semantic errors leading to failed tests and crashing errors that terminate execution abruptly. Further, they reflect two important, widely-used modalities for Python programming: non-interactive scripts and interactive computational notebooks. ChatDBG supports debugging in both settings.
Programs were prepared by removing them from their automatic grading harness and replacing failed unit tests with assert statements that generate exceptions. We ran each program ten times under the five configurations in Table IV: Default Stack includes standard stack traces, as generated by ipdb [23], with 5 lines of code per frame in the initial prompt, but it does not support the LLM taking the wheel. Enriched Stack generates enriched stacks with ten lines of code per frame, and +Take the Wheel additionally permits ChatDBG to run debugger commands. These three configurations all use why? as the initial user text. +Targeted Question asks a question specific to the failure. For semantic errors, which validate the values stored in variables, these questions describe what those values should be or what they are intended to represent. For crashes, the questions relate the crash to expected behavior, as in the following; we designed our questions to be “neutral” and not hint at the root cause.
c3 (Crash)
Why am I not reading the CSV file correctly?
s11 (Crash)
Why am I not able to sample 100 rows?
c1 (Semantic)
Why am I not getting 3?
s1 (Semantic)
bill_length_mean_by_species should be a table of the mean bill lengths of each species in our data set. Why isn’t it?
The final +Dialog configuration is the same as +Targeted Question but extends the chat with a second query. All trials use the same follow up: Continue to explain your reasoning and give me a fix to make it work as I describe. Context-specific follow-ups work better, but we opted for consistency.
ChatDBG used the gpt-4-1106-preview model for these experiments. Under +Targeted Question, the first prompt and response led to, on average, a chat of about 10,000 tokens (7,500 words), a cost of about \$$0.120.12$ USD under OpenAI’s current pricing model [24], and a completion time of about 25 seconds. Subsequent steps in extended debugging dialogs incurred comparable costs. Time was highly variable and dominated by the performance of OpenAI’s service. These characteristics will be different for other platforms and models.
RQ1: Is ChatDBG effective at diagnosing and fixing bugs in Python?
Each response was manually examined and deemed a success if it included an accurate explanation of the error and an actionable fix. That fix could be either code or a prose description in which all necessary details were made explicit. Figure 6 shows the success rate under each configuration.
RQ1 Summary: Even with just the simple question why?, ChatDBG was successful 57% of the time. With questions specialized to the target’s particular error, that number jumps to 67%, and with an additional dialog step ChatDBG succeeded in identiyfing and fixing the defect in 85% of the trials.
Configuration Stack Take the Initial Ask a
Trace Wheel Prompt Follow-up
Default Stack standard why?
Enriched Stack enriched why?
+Take the Wheel enriched why?
+Targeted Question enriched specialized
+Dialog enriched specialized
TABLE IV: Configurations used in the Python experiments.
Figure 6: Overall ChatDBG success rate for each configuration. ChatDBG innovations and user-provided context gradually increase effectiveness.
Figure 7: Success rate for ChatDBG for each program and configuration. Vertical lines show the mean.
Program LoC Error Type Root Cause   Proximate Cause Fix Root Cause Fix
BC [25] 17.0k Buffer overflow Input from data file printed to a fixed-size buffer Truncate on copy Use dynamic size
GZIP [25] 8.2k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
NCOM [25] 1.9k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
PEG [26] 14.7k Null dereference Invalid input produces corrupted data structure Check if not null Warn/exit
POLY [25] 0.7k Buffer overflow Command line argument is unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
TIFF [26] 58.9k Division by zero Combination of command line options leads to a division by zero Override option values Warn/exit when invalid
YAML1 [26] 8.7k Stack overflow Long sequences of { in the input leads to deep recursion Use iterative method Guard recursion depth
YAML2 [26] 8.7k Assertion failure Specific input causes a peek request for non-existent “next” token Replace assert Check before peeking
TABLE V: Bugs in unmanaged C/C++ code, and our criteria for fixing the proximate cause or the root cause of each.
RQ2: Which components of ChatDBG contribute to its effectiveness?
Figure 7 presents the success rates for each program under each configuration. The Enriched Stack plots demonstrate that enriched stacks provide some benefit, particuarly for crashes in which the stack contains sufficient information to diagnose the problem, but they alone do not provide much improvement for many semantic errors in which the relevant computation steps complete before failure. However, enriched stacks coupled with letting the LLM taking the wheel led to significant improvement in the success rate for both crashing and semantic bugs, as shown in the +Take the Wheel plots.
The LLM most heavily used the info, slice (for notebooks), and p (print) debugging commands. That is not surprising, given that they often provide the most direct insight into the execution state and code. The slice command was critically important for notebooks. Without it, success rates rarely improved when the LLM took the wheel.
The +Targeted Question configuration demonstrates the impact of providing even the most modest details about expected behavior. When the LLM is asked to continue its reasoning in +Dialog, ChatDBG’s success rate improves despite the follow-up prompt providing no feedback on the contents or quality of the first response. This phenom indicates that constraints on the underlying LLM’s response lengths may prevent it from conducting the amount of reasoning necessary to successfully develop a fix in a single step. The success rates for +Targeted Question and +Dialog demonstrate the importance of continued dialogs and user input. We expect those features to be even more important to ChatDBG’s success when diagnosing bugs in more complex programs.
The LLM also demonstrated its background knowledge with the responses including, for example, details of Python idioms and libraries, the definition of h-index [27], and the implementation and limitations of various statistical techniques.
Failures were generally due to the LLM not always recognizing or discovering key aspects of a program’s behavior. In some cases, it was on the right track but did not converge on an actionable fix. In others, it suggested changes that would introduce other bugs. It also occasionally made mistakes, such as conflating proportions and percentages or failing to handle unusual corner cases. All of these could be mitigated by feedback from the user in subsequent follow ups.
RQ2 Summary: While all features of ChatDBG contribute to its success, the technical innovations enabling it to take the wheel are critical. The most sophisticated configurations show that user-provided contextual information about behavior and engaging in multi-step dialogs are particularly good ways to improve its effectiveness.
V-B C and C++
Programs in unmanaged languages such as C and C++ are vulnerable to memory safety errors. These memory errors can also hinder the debugging process: the crash may not occur immediately at the memory violation but instead much later on, and the crash may cause corruption of the stack and/or heap, making it challenging to recover any useful information.
Table V summarizes the programs from the BugBench [25] and BugsC++ [26] suites used to evaluate ChatDBG’s effectiveness at debugging unmanaged code. Programs used in this evaluation are all real-world applications with concrete known bugs. The four BugBench programs were selected as the only ones we could retrieve, build and reproduce on our system. The BugsC++ suite does not include the original crash-causing inputs. However, it provides links to the original bug report, CVE identifier, and/or exploit-fixing patch, from which we manually retrieve crash reproduction information. We randomly selected and reproduced four bugs from the “memory error” category.
Some of the studied programs do not crash at run time. We employed AddressSanitizer [28] to force a crash at the moment a memory violations occured to trigger those defects. AddressSanitizer is already capable of reporting some information about the crash when it happens. However, this information is often very dense, and typically points at the symptom of the bug, not its root cause. We did not include that information in the initial prompt.
RQ3: Is ChatDBG effective at diagnosing and fixing bugs in unmanaged code (C/C++)?
We run our C/C++ experiments on an an x86 Ubuntu 22.04 server. We use Clang and LLDB 17.0.6 to compile and debug, using flags -g -Og -fno-omit-frame-pointer. ChatDBG used OpenAI’s gpt-4-1106-preview model. Each program was run ten times using queries of the form I am debugging cpp-peglib. Provide the root cause of this crash, for PEG, followed by a request to include code in the response. Average time (27 seconds) and cost (\$$0.060.06$ USD) were comparable to Python.
We manually examined each response to determine if ChatDBG successfully provided an actionable code fix for the proximate cause of the crash or for the underlying root cause. We used the critera outlined in Table V. While fixing root causes is the ultimate goal, fixing proximate causes can still be beneficial as fixing crashes enables further debugging steps.
Figure 8: ChatDBG success rate at fixing the proximate or root cause in C/C++ programs. ChatDBG successfully identified and fixed the root cause 36% of the time and the proximate cause an additional 55% of the time.
Figure 8 presents ChatDBG’s ability to suggest a fix for either the proximate or root cause of the bug. Generally, ChatDBG is excellent at diagnosing and explaining the reason for the crash, which in itself may be useful to programmers. For BC, GZIP, NCOM, and POLY, ChatDBG tends to suggest replacing the strcpy or sprintf call with their respective strncpy and snprintf counterparts to prevent buffer overflows. While correct, this change truncates the input silently. Validation or other measures should be added to obtain a robust fix. The root cause in BC is inside code generated from a YACC file. The clangd language server does not handle this case in a way that would let ChatDBG answer the LLM’s definition requests properly.
In the case of PEG, ChatDBG correctly identifies which pointer is null but typically suggests ignoring it instead of failing immediately. This is similar to YAML2, where ChatDBG recommends replacing the assertion with a check inside a function rather than recommending that the client check that the function’s preconditions are met prior to the call. ChatDBG has a relatively high root cause fix rate for YAML1 and TIFF. It often correctly suggests fixes to limit recursion depth (YAML1) and to validate input parameters (TIFF).
RQ3 Summary: ChatDBG was successful in virtually all of our trials in diagnosing and explaining the cause of the crash. It was also capable of providing relevant, actionable fixes: 36% of its suggestions addressed the root cause of the bug, while another 55% corrected the proximate cause.
V-C Threats to Validity
This paper evaluates ChatDBG on two suites of code. The primary suite is a collection of unpublished student labs that may not be entirely representative of code written by, for example, experienced programmers. The second suite consists of real C/C++ applications and bugs drawn from the BugBench and BugsC++ suites. Unlike the Python suite, the C/C++ source code and the bug fixes for these programs are available on Github, which may lead to data leakage affecting the C/C++ study if those repositories were part of the training set for the LLMs we used. While the C/C++ suite consists of real-world applications, most of the errors are memory errors. Other types of errors, such as assertion failures, concurrency errors, or other logical errors, may lead to different results.
ChatDBG depends on an LLM to analyze and drive exploration of state, and like all systems based on LLMs today, its performance is affected by prompt engineering. It is possible that ChatDBG’s prompts are overfit to the specific GPT-4 models we employed; this threat is somewhat mitigated by the fact that ChatDBG was originally developed using a different model (GPT-3.5-turbo). LLMs are also inherently stochastic, and it is possible to obtain unusually good results by chance. To mitigate this threat, our evaluation runs ChatDBG on each test program at least ten times.
Our evaluation depends on a manual and subjective evaluation of whether ChatDBG’s explanation of a bug and its proposed fix are satisfactory. We mitigated the risks of subjective evaluation by using precisely-defined criteria decided upon in advance. Python fixes were deemed successful if the resulting code met the correctness requirements outlined in the assignment. C/C++ fixes were deemed successful at fixing the proximate or root cause using the criteria in Table V. Fixes described in prose were permitted, provided that the details of all necessary changes to the code were made explicit.
VI Conclusion and Future Work
This paper presents ChatDBG, the first AI-based debugging assistant. Our evaluation shows that engaging in a debugging dialog with ChatDBG can significantly assist in identifying root causes of errors and developing correct fixes.
We see several avenues for future work. Incorporating an existing fault localization approach into ChatDBG, rather than relying solely on the LLM’s ability to explore the program’s source code and state, could potentially increase its effectiveness and efficiency by allowing the LLM to focus its attention on suspicious and possibly problematic files, functions, or lines of source code. Similarly, incorporating delta debugging [29] could increase the effectiveness of ChatDBG by limiting the amount of input for an LLM and providing failure-inducing events as guidance. Finally, integrating ChatDBG with a time-travel debugger would expand its reach to exploring program state over time, letting it answer queries that cannot be answered given the current program state.
ChatDBG is available on GitHub at github.com/plasma-umass/ChatDBG.

### v2

@@
 ChatDBG: An AI-Powered Debugging Assistant
-Abstract
-This paper presents ChatDBG, the first AI-powered debugging assistant.
+Abstract.
+Debugging is a critical but challenging task for programmers.
+This paper proposes ChatDBG, an AI-powered debugging assistant.
 ChatDBG integrates large language models (LLMs) to significantly enhance the capabilities and user-friendliness of conventional debuggers.
-ChatDBG lets programmers engage in a collaborative dialogue with the debugger, allowing them to pose complex questions about program state, perform root cause analysis for crashes or assertion failures, and explore open-ended queries like why is x null?.
-To handle these queries, ChatDBG grants the LLM autonomy to take the wheel and drive debugging by issuing commands to navigate through stacks and inspect program state; it then reports its findings and yields back control to the programmer.
-Our ChatDBG prototype integrates with standard debuggers including LLDB, GDB, and WinDBG for native code and Pdb for Python.
+ChatDBG lets programmers engage in a collaborative dialogue with the debugger, allowing them to pose complex questions about program state, perform root cause analysis for crashes or assertion failures, and explore open-ended queries like ‘why is x null?’.
+To handle these queries, ChatDBG grants the LLM autonomy to take the wheel: it can act as an independent agent capable of querying and controlling the debugger to navigate through stacks and inspect program state.
+It then reports its findings and yields back control to the programmer.
+Our ChatDBG prototype integrates with standard debuggers including LLDB and GDB for native code and Pdb for Python.
 Our evaluation across a diverse set of code, including C/C++ code with known bugs and a suite of Python code including standalone scripts and Jupyter notebooks, demonstrates that ChatDBG can successfully analyze root causes, explain bugs, and generate accurate fixes for a wide range of real-world errors.
 For the Python programs, a single query led to an actionable bug fix 67% of the time; one additional follow-up query increased the success rate to 85%.
-ChatDBG has seen rapid uptake; it has already been downloaded nearly 30,000 times.
-11footnotetext: Equal contribution.footnotetext: Work done at the University of Massachusetts Amherst.
-I Introduction
+ChatDBG has seen rapid uptake; it has already been downloaded roughly 50,000 times.
+Debugging, Artificial Intelligence, Software Engineering
+$c$s: Computing methodologies Artificial intelligenceccs: Software and its engineering Software testing and debuggingconference: Under Submission; 2025; Someplace, Somewhere
+1. Introduction
 Debuggers help programmers identify and fix bugs by letting them investigate program state and navigate program execution.
-Debuggers for mainstream languages, including GDB [1], LLDB [2], and WinDBG (for C, C++, and Rust), jdb (Java), Pdb (Python), and the Chrome or Firefox debuggers (for JavaScript), generally provide the same functionality.
+Debuggers for mainstream languages, including GDB (Stallman et al., 2011) and LLDB (LLDB, 2010) (for C, C++, and Rust), JDB (for Java), Pdb (for Python), and the Chrome or Firefox debuggers (for JavaScript), generally provide the same functionality.
 In particular, most debuggers support observing program execution via tracing and reporting when a program reaches a given line or function of source code; interrupting execution and returning control to the debugger when the program reaches a given line or function via breakpoints, when a particular condition is true via conditional breakpoints, or when a variable changes via watchpoints (a.k.a.
-data breakpoints); inspecting local variables, globals, heap objects, and backtraces of the call stack; resuming program execution line-by-line (single-step) or at the granularity of function calls; and in some debuggers, stepping backward through execution via reverse debugging, also known as time-travel or omniscient debugging.
+data breakpoints); inspecting local variables, globals, heap objects, and backtraces of the call stack; and resuming program execution line-by-line (single-step) or at the granularity of function calls.
 Debuggers can be helpful, but finding and fixing software defects remains a deeply challenging and time-consuming task.
@@
 ChatDBG integrates into and significantly extends the functionality of standard debuggers.
-ChatDBG builds on the insight that large language models (LLMs), such as OpenAI’s GPT-4 [3], enable a debugger to leverage insights and intuition from many thousands of programs as well as the vast real-world knowledge embedded in LLMs.
-A debugger integrated with ChatDBG continues to provide its full range of functionality, but also lets programmers engage in debugging dialogs where they can ask high-level questions like why is x null here?
-or why isn’t this value what I expected?.
-The question can be as simple as why?
-if a program has crashed or failed an assertion.
-ChatDBG then orchestrates a conversation with the LLM.
-A key novelty of ChatDBG is that it grants autonomy to the LLM to “take the wheel” while answering the programmer’s queries.
-Specifically, the LLM issues “function calls” [4] to run commands in the underlying debugger to investigate program state, execute code, or obtain source code.
+ChatDBG builds on the insight that large language models (LLMs), such as OpenAI’s GPT-4 (OpenAI, 2023b), enable a debugger to leverage insights and intuition from the vast real-world knowledge embedded in LLMs.
+A debugger integrated with ChatDBG continues to provide its full range of functionality but also lets programmers engage in debugging dialogs where they can ask high-level questions like ‘why is x null here?’ or ‘why isn’t this value what I expected?’.
+The question can be as simple as ‘why?’ if a program has crashed or failed an assertion.
+To answer such queries, ChatDBG orchestrates a conversation with an LLM.
+A key novelty of ChatDBG is that it grants autonomy to the LLM to “take the wheel” and act as an independent agent while answering the programmer’s queries.
+Specifically, the LLM issues “function calls” (OpenAI, 2023a) to run commands in the underlying debugger to investigate program state, execute code, or obtain source code.
 The results of those calls are sent back to the LLM for use while constructing its response.
 After answering a query, control is returned to the programmer, who may then enter additional commands or chat messages.
-Our prototype of ChatDBG integrates into four widely used debuggers: GDB, LLDB, WinDBG, and Pdb.
+Support for the LLM’s agentic control of the debugger is a key technical contribution.
+Our prototype of ChatDBG integrates into three widely used debuggers: GDB, LLDB, and Pdb.
 Our evaluation presents a range of case studies demonstrating that ChatDBG improves significantly on existing debuggers.
-On a suite of unpublished Python scripts and Jupyter notebooks written by undergraduate students, one or two queries is sufficient for ChatDBG to properly diagnose and fix defects 87% of the time, typically at a cost of under \$0.20 USD.
+On a suite of unpublished Python scripts and Jupyter notebooks written by undergraduate students, one or two queries is sufficient for ChatDBG to properly diagnose and fix defects 87% of the time, typically at a cost well under \$0.20 USD.
 ChatDBG is also effective at identifying causes and providing fixes for a range of real-world bugs in C/C++ code.
-This paper makes the following contributions: it introduces ChatDBG, the first AI-powered debugger assistant; it describes the implementation of our ChatDBG prototype; and it introduces the “take the wheel” approach to integrating large language models with developer tools.
-The paper also presents an evaluation of ChatDBG that demonstrates its significant advantages over existing debugger functionality.
+This paper makes the following contributions.
+  • It introduces ChatDBG, the first AI-powered debugger assistant.
+  • It describes the implementation of our ChatDBG prototype.
+  • It introduces the “take the wheel” approach to integrating agentic reasoning in large language models with developer tools.
+  • It presents an evaluation of ChatDBG that demonstrates its significant advantages over existing debugger functionality.
+ChatDBG is broadly applicable to many domains and programming languages, and we expect it to be particularly useful for novice programmers, who often lack the experience to effectively use debuggers.
+ChatDBG is also useful for experienced programmers, who can augment debugging sessions with ChatDBG’s reasoning capabilities in a natural, conversational, and interactive way.
 This paper is organized as follows.
-Section II illustrates the use of ChatDBG to debug a program.
-Section III describes key related work.
-Section IV describes ChatDBG’s implementation.
-Section V presents our evaluation, and Section VI concludes with several promising directions for future work.
+Section 2 illustrates the use of ChatDBG to debug a program.
+Section 3 describes key related work.
+Section 4 describes ChatDBG’s implementation.
+Section 5 presents our evaluation.
+Section 6 outlines directions for future work, and Section 7 concludes.
 Source code for bootstrap.py
-from datascience import *
-from ds101 import *
-def make_marble_sample():
-    table = Table().read_table(’marble-sample.csv’)
-    return table.column(’color’)
-def proportion_blue(sample):
-    return sample
-def resampled_stats(observed_marbles, num_trials):
-    stats = bootstrap_statistic(observed_marbles,
-                                proportion_blue,
-                                num_trials)
-    assert len(stats) == num_trials
-    return stats
-observed_marbles = make_marble_sample()
-stats = resampled_stats(observed_marbles, 5)
-assert np.isclose(np.mean(stats), 0.7)
-Figure 1: An example program containing several bugs.
+[…]
+[fontsize=,xleftmargin=20pt,linenos]pythonfigs/bootstrap.py
+Figure 1.
+An example program containing several bugs.
 It is supposed to create an array of marble colors, compute the proportions of blue marbles in resamples of that array, and assert that their mean is about 0.7, the proportion for the array.
@@
 Entering post mortem debugging Running ’cont’ or ’step’ will restart the program (ChatDBG) p num_trials 5 (ChatDBG) Why doesn’t stats have 5 elements?
-(ChatDBG) p len(stats) 150 (ChatDBG) info bootstrap_statistic bootstrap_statistic(observed_sample,                     compute_statistic, num_trials)    Creates num_trials resamples of the initial    sample.
+(ChatDBG) p len(stats) 150
+(ChatDBG) info bootstrap_statistic bootstrap_statistic(observed_sample, compute_statistic, num_trials)     Creates num_trials resamples of the initial sample.
 Returns an array of the     provided statistic for those samples.
@@
 * compute_statistic: a function that takes a sample as an array and returns       the statistic for that sample.
-* num_trials: the number of bootstrap samples                  to create.You MUST assume that ‘bootstrap_statistic‘ isspecified and implemented correctly.
-(ChatDBG) p proportion_blue <function proportion_blue at 0x1050b00e0> (ChatDBG) info proportion_blue   8     def proportion_blue(sample):   9         return sample From the information we have, the stats variable in resampled_stats is supposed to have 5 elements, but it actually has 150 elements… The Python debugger output shows a reference to a variable or function called proportion_blue, which is used as the compute_statistic parameter for bootstrap_statistic.
+* num_trials: the number of bootstrap samples to create.
+You MUST assume that ‘bootstrap_statistic‘ is specified and implemented correctly.
+(ChatDBG) p proportion_blue <function proportion_blue at 0x1050b00e0>
+(ChatDBG) info proportion_blue   8     def proportion_blue(sample):   9         return sample
+{adjustwidth}
+.5em1em From the information we have, the stats variable in resampled_stats is supposed to have 5 elements, but it actually has 150 elements…
+The Python debugger output shows a reference to a variable or function called proportion_blue, which is used as the compute_statistic parameter for bootstrap_statistic.
 Upon looking at the proportion_blue function’s definition, it is found that the function simply returns the input sample without computing a statistical value.
@@
 Recommendation
-To fix this issue, the proportion_blue function must be corrected to perform the intended computation of returning the proportion of blue marbles in the sample: def proportion_blue(sample):    return np.mean(sample == 'B')
-Figure 2: A debugging session with ChatDBG.
+To fix this issue, the proportion_blue function must be corrected to perform the intended computation of returning the proportion of blue marbles in the sample:
+def proportion_blue(sample):    return np.mean(sample == 'B')
+Figure 2.
+A debugging session with ChatDBG.
 For brevity, we elide repetitive or unimportant parts.
@@
 Commands issued by ChatDBG when it takes the wheel are highlighted in yellow, and its response is highlighted in blue.
-II Overview
+2. Overview
 This section illustrates ChatDBG’s features and ability to assist in debugging the program in Figure 1.
@@
 In this case, the statistic is proportion_blue, the proportion of blue marbles.
-Given a sufficiently large number of trials, the mean of the resamples’ statistics should be close to 0.7, the proportion of blue marbles in the original sample [5].
+Given a sufficiently large number of trials, the mean of the resamples’ statistics should be close to 0.7, the proportion of blue marbles in the original sample (Efron, 1979).
 The program fails the assertion in resampled_stats, and Figure 2 illustrates a debugging session.
@@
 While constructing the answer (in blue), the LLM takes the wheel and directly issues debugger commands (yellow).
-These include standard Pdb commands and a ChatDBG-specific info command for accessing the source code and docstrings for any user-written code, and the docstrings for library code (which we assume is correct and not the root cause of any error).
+These include standard Pdb commands and a ChatDBG-specific info command for accessing the source code and docstrings for any user-written code, as well as the docstrings for library code (which we assume is correct and not the root cause of errors).
 \$ chatdbg bootstrap.py Traceback (most recent call last):   File "./bootstrap.py", line 21, in <module>     assert np.isclose(np.mean(stats), 0.7)            ^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^^ AssertionError ...
 (ChatDBG) Why is the mean of stats not close to 0.7?
-(ChatDBG) p np.mean(stats) 0.7466666666666667 (ChatDBG) info resampled_stats ...
+(ChatDBG) p np.mean(stats) 0.7466666666666667
+{adjustwidth}
+.5em1em …
+(ChatDBG) info resampled_stats ...
 (ChatDBG) info make_marble_sample ....
 (ChatDBG) info bootstrap_statistic ...
+{adjustwidth}
+.5em1em
 Recommendation
-1.
+  1. 1.
     Check the CSV file marble-sample.csv to verify the proportion of blue to red marbles and ensure it is close to 70% blue marbles.
-2.
-Review the implementation of the proportion_blue function… 3.
+  2. 2.
+    Review the implementation of the proportion_blue function…
+  3. 3.
     Verify that the proportion_blue function … is available in the current scope or imported correctly.
 (ChatDBG) p proportion_blue(observed_marbles) 0.7 (ChatDBG) That value is correct.
-(ChatDBG) p stats array([0.66666667, 0.8       , 0.63333333,        0.76666667, 0.86666667]) At this point, the most straightforward explanation for the error is that the number of trials is not enough to accurately reflect the expected mean due to random sampling effects.
+{adjustwidth}
+.5em1em
+(ChatDBG) p stats array([0.66666667, 0.8       , 0.63333333,        0.76666667, 0.86666667])
+{adjustwidth}
+.5em1em
+At this point, the most straightforward explanation for the error is that the number of trials is not enough to accurately reflect the expected mean due to random sampling effects.
 Increasing the number of trials would likely result in an average statistic that is closer to the expected value of 0.7, following the Law of Large Numbers.
@@
 Increase the number of bootstrap trials significantly (e.g., from 5 to something like 1000 or more) to reduce the variability of the mean…
-Figure 3: A debugging session demonstrating how ChatDBG incorporates real-world knowledge.
+Figure 3.
+A debugging session demonstrating how ChatDBG incorporates real-world knowledge.
 After correcting proportion_blue in Figure 1, the program fails on line 21 because the mean proportion of blue marbles in the resamples is not the expected value.
@@
 A powerful aspect of ChatDBG is its ability to exploit real-world knowledge in its analyses (here, the fact that bootstrapping depends on large numbers of resamples) without specific instruction or user intervention.
-III Related Work
-System & Date Single Stack Break- Cond. Source Trace Display Eval. Watch- Reverse Explain Propose Open
-Introduced Step Nav. points " Level State Code points Bugs Fixes Queries
-DDT [6], 1961
-EXDAMS [7], 1969
-Mesa [8], 1979 ✓*
-DBX [9], 1981 1990
-GDB [10], 1986 1991 2009
+3. Related Work
+System and Date Single Step Stack Navigation Breakpoints (BPs) Conditional BPs Source Level Trace Display State Eval. Code Watchpoints Explain Bugs Propose Fixes
+Open Queries
+DDT (Kotok, 1961), 1961
+EXDAMS (Balzer, 1969), 1969
+Mesa (Xerox, Systems Development Department, 1979), 1979 ✓*
+Dbx (Linton, 1990), 1981 1990
+GDB (Stallman et al., 2011), 1986 1991
 Pdb, 1992
-WinDBG, ca. 1997 2017
 LLDB, 2010
 ChatDBG, 2023
-TABLE I: Debugger features and their dates of introduction.
+Table 1.
+Debugger features and their dates of introduction.
 Most key features have been around for decades.
-By integrating into modern debuggers (GDB, LLDB, Pdb and WinDBG), ChatDBG inherits all of their features while significantly extending them with functionality to explain bugs and their root causes, propose fixes, and answer arbitrary natural-language queries over program state.
-(An asterisk or year in italics means the feature is limited in functionality, performance, or depends on specific hardware support; for example, WinDBG’s reverse debugging only works on specific Intel chips.)
-Table I presents an overview of previous interactive debuggers, together with their features and date of introduction.
-The first interactive debugger, DDT, introduced breakpoints, single-stepping, and stack navigation in 1961 [6].
-By 1979, the Mesa debugger had most key features of modern debuggers, including source-level debugging, conditional breakpoints, tracing, and the ability to display runtime state and evaluate code [8].
-Arbitrary conditional breakpoints date back at least to 1990 with DBX [9].
-Watchpoints were introduced by 1991 and have been in GDB since version 4.0.1 [10].
-Reverse debugging it was first implemented in EXDAMS in 1969 [7] is present in widely-used debuggers like GDB (in 2009) [11], and WinDBG (in 2017); the rr debugger, built on top of GDB, also supports reverse debugging on Linux platforms [12].
-ChatDBG, by integrating into standard debuggers, inherits and extends their functionality.
-Ko and Myers present Whyline, an interactive, trace-based debugger that lets programmers select from a range of queries and identifies (via static and dynamic analysis) a timeline that answers the query [13].
+By integrating into modern debuggers (GDB, LLDB, and Pdb), ChatDBG inherits all of their features while significantly extending them with functionality to explain bugs and their root causes, propose fixes, and answer arbitrary natural-language queries over program state.
+(An asterisk or year in italics means the feature is limited in functionality, performance, or depends on specific hardware support.)
+Table 1 presents an overview of previous interactive debuggers, together with their features.
+The first interactive debugger, DDT, introduced breakpoints, single-stepping, and stack navigation in 1961 (Kotok, 1961).
+By 1979, the Mesa debugger had most key features of modern debuggers, including source-level debugging, conditional breakpoints, tracing, and the ability to display run-time state and evaluate code (Xerox, Systems Development Department, 1979).
+Arbitrary conditional breakpoints date back at least to 1990 with Dbx (Linton, 1990).
+Watchpoints were introduced by 1991 and have been in GDB since version 3.93.
+In other work, Ko and Myers present Whyline, an interactive, trace-based debugger that lets programmers select from a range of queries and identifies (via static and dynamic analysis) a timeline that answers the query (Ko and Myers, 2010).
 Programmers can only select from those queries presented by Whyline as options.
@@
 Whyline’s use of traces gives it the ability to answer questions that might not be straightforward to answer with the current program state but limits its applicability to relatively short-lived executions.
-The goal of program slicing, introduced by Weiser in 1981 [14], is to produce a shorter version of a program limited to the source code that could have led to an error.
+The goal of program slicing, introduced by Weiser in 1981 (Weiser, 1981), is to produce a shorter version of a program limited to the source code that could have led to an error.
 Program slicing has been extensively studied; Weiser’s paper has been cited over 5,000 times to date.
-As Section IV-G describes, ChatDBG performs backwards slicing to collect code spread across code cells to facilitate debugging of Jupyter notebooks.
-Fault localization seeks to identify the likely location of the root cause of a defect; Wong et al.
-present a survey with over 400 citations [15].
-Unlike previous work, ChatDBG performs fault localization by leveraging LLM-based examination and LLM-driven exploration of source code and program state.
-Automated program repair is another highly active area of software engineering research [16]; its goal is to generate source-level program patches that prevent a program from failing.
-Unlike past work, ChatDBG performs best-effort automated program repair—leveraging its integration within a debugger—via LLM-based examination and LLM-driven exploration of program source and state.
-IV Implementation
-[…]
-1:history = “”
-2:loop
-3:    line = Input()
-4:    if IsDebuggerCommand(line) then
-[…]
-6:        Print(output)
-7:        history = history + (line + $\rightarrow$” + output)
-8:    else
-[…]
-10:             prompt = MakePrompt(Instructions(),
-11:                                    EnrichedStack(), Inputs(),
-12:                                    Error(), history, line)
-13:         else
-14:             prompt = MakePrompt(history, line)
-[…]
-16:         history = “”
-17:         while ResponsePending() do
-18:             match Receive()
-[…]
-20:                      output = DoCommand(cmd)
-21:                      Print (cmd + $\rightarrow$” + output)
-22:                      Send(output)
-[…]
-Figure 4: ChatDBG architecture and top-level command processing loop.
-IV-A Using ChatDBG: Preliminaries
+As Section 4.7 describes, ChatDBG performs backwards slicing to collect code spread across code cells to facilitate debugging of Jupyter notebooks.
+Fault localization seeks to identify the likely location of a defect’s root cause.
+Several prior studies have investigated the use of machine learning and LLMs for fault localization.
+Some of the studied techniques apply machine learning to source code features, coverage data, or other static code features to predict faulty lines of code, but they do not utilize dynamic state and run-time information.
+DeepFL (Li et al., 2019), Grace (Lou et al., 2021) and DeepRL4FL (Li et al., 2021) are examples of such systems.
+Similarly, LLMAO (Yang et al., 2024) employs LLMs to provide suspiciousness scores for each line of code in a given program, but only provides access to the source code.
+AutoFL (Kang et al., 2024) also utilizes an LLM and enables it to statically retrieve source code and coverage information about the program via function calls.
+However, the system requires a failing test case as input and does not employ run-time state information.
+ChatDBG improves upon these systems by providing an LLM with access to run-time program state and the ability to take control of the underlying debugger.
+Both of these features enhance the LLM’s ability to provide more accurate and informative feedback to the user.
+We also note that other fault localization techniques can be used in tandem with ChatDBG to improve results, as suggested by Section 5.1’s utilization of backwards slicing to identify code relevant to a bug in Python notebooks and Section 5.2’s utilization of AddressSanitizer (Serebryany et al., 2012) to provide a better starting point for diagnosing and fixing memory errors in native code.
+Automated program repair is another active area of software engineering research (Goues et al., 2019).
+Systems for automatic program repair attempt to generate source-level program patches that prevent a program from failing.
+ChatDBG performs best-effort automated program repair by requesting that the LLM propose code fixes as part of its response, ultimately letting the programmer drive code changes using these suggestions.
+Previous research has shown that automated program repair hints can provide significant help in the debugging process and suggests that the benefits of correct advice outweigh the risk of deceptive ones (Eladawy et al., 2024).
+4. Implementation
+4.1. Using ChatDBG: Preliminaries
 ChatDBG integrates with existing debuggers as either a plug-in or a direct extension.
@@
 Since ChatDBG leverages OpenAI’s LLMs, the user must also set an environment variable to a valid OpenAI API key within their system’s configuration settings.
+4.2. Debugging a Target Program
+For Python, debugging with ChatDBG begins by running chatdbg on the target program.
+No special preparation of the target is needed; Python’s managed run time ensures that debugging information and source code is always available.
+Debugging is supported in IPython interactive sessions or Jupyter notebooks via the standard command-line flag --pdb or the Jupyter magic command %pdb, respectively.
+Control drops into the debugger when an exception occurs.
+For C and C++, debugging begins by running lldb on the target program.
+The target program must be an unstripped executable generated with the -g compiler flag, which ensures the availability of DWARF debug information that describes the memory layout and maps the program’s machine code back to the original source code.
+That information is essential for the effective debugging of unmanaged code.
+ChatDBG also handles native code generated for other languages but may require additional steps.
+For example, to debug a Rust target program, the Cargo.toml file must list ChatDBG as a dependency and the main function must be annotated with #[chatdbg::main] to ensure that error messages are visible to ChatDBG through a log file.
+[…]
+Figure 4. ChatDBG architecture and command processing algorithm.
+4.3. ChatDBG Architecture Overview
+ChatDBG orchestrates communication between the user, the debugger, and the LLM, as shown in the architecture diagram in Figure 4.
+The operations in the command loop pseudocode map naturally onto debugger APIs and onto LLM APIs supporting completion and function calls (OpenAI, 2023a).
+ChatDBG currently utilizes OpenAI’s API (OpenAI, 2020) and GPT-4 models.
+We provide a brief overview of the ChatDBG architecture and then elaborate on the most salient technical innovations below.
+[…]
+[…]
+{justify}
 Instructions:
-  You are a debugging assistant.
-You will be given a Python stack trace for an error and answer questions related to the root cause of the error.
+  • You are a debugging assistant. You will be given a Python stack trace for an error and answer questions related to the root cause of the error.
   • Call the debug function to run Pdb debugger commands on the stopped program.
@@
 Your answer may be as long as necessary.
-End your answer with a section titled “Recommendation” that contains one of: a fix if you have identified the root cause a numbered list of 1-3 suggestions for how to continue debugging if you have not
+  • End your answer with a section titled “Recommendation” that contains one of:
+    • a fix if you have identified the root cause
+    • a numbered list of 1-3 suggestions for how to continue debugging if you have not
 Enriched Stack Trace:
-  The program has this stack trace: [...
-skipping 4 hidden frame(s)] ./bootstrap.py(19)<module>()      15      16     assert len(stats) == num_trials      17      18 observed_marbles = make_marble_bag() ---> 19 resampled_stats(observed_marbles, 5)    Global variables:      observed_marbles: ndarray = array([’R’, R’, R’,        ..., B’, B’, B’], dtype=’<U1’) > ./bootstrap.py(16)resampled_stats()      14                                 num_trials)      15 ---> 16     assert len(stats) == num_trials      17      18 observed_marbles = make_marble_bag()    Variables in this frame:      num_trials: int = 5      observed_marbles: ndarray = array([’R’, R’, R’,        ..., B’, B’, B’], dtype=’<U1’)      stats: ndarray = array([’B’, R’, B’, ..., R’,        B’, R’], dtype=’<U32’)
+The program has this stack trace:
+[...
+skipping 4 hidden frame(s)]./bootstrap.py(19)<module>()     15     16     assert len(stats) == num_trials     17     18 observed_marbles = make_marble_bag()---> 19 resampled_stats(observed_marbles, 5)   Global variables:     observed_marbles: ndarray = array([’R’, ’R’, ’R’, ..., ’B’,                                        ’B’, ’B’], dtype=’<U1’)> ./bootstrap.py(16)resampled_stats()     14                                 num_trials)     15---> 16     assert len(stats) == num_trials     17     18 observed_marbles = make_marble_bag()   Variables in this frame:     num_trials: int = 5     observed_marbles: ndarray = array([’R’, ’R’, ’R’, ...,                                        ’B’, ’B’, ’B’], dtype=’<U1’)     stats: ndarray = array([’B’, ’R’, ’B’, ..., ’R’, ’B’,                             ’R’], dtype=’<U32’)
 Error:
-  The program encountered the following error: AssertionError The code assert len(stats) == num_trials is correct and must not be changed.
+The program encountered the following error:
+AssertionErrorThe code assert len(stats) == num_trials is correct and must not be changed.
 History:
-  This is the history of some pdb commands I ran and the results:   (ChatDBG) p num_trials   5
+This is the history of some pdb commands I ran and the results:
+(ChatDBG) p num_trials5
 User Text:
 Why doesn’t stats have 5 elements?
-Figure 5: The initial prompt for the debugging session in Figure 2. For brevity, the enriched stack includes only five lines of source in each frame, rather than the default of 10.
-IV-B Debugging a Target Program
-For Python, debugging with ChatDBG begins by running chatdbg on the target program.
-No special preparation of the target is needed; Python’s managed run time ensures that debugging information and source code is always available.
-Debugging is supported in IPython interactive sessions or Jupyter notebooks via the standard command-line flag --pdb or the Jupyter magic command %pdb, respectively.
-Control drops into the debugger when an exception occurs.
-For C and C++, debugging begins by running lldb on the target program.
-The target program must be an unstripped executable generated with the -g compiler flag, which ensures the availability of DWARF debug information that describes the memory layout and maps the program’s machine code back to the original source code.
-That information is essential for the effective debugging of unmanaged code.
-ChatDBG also handles native code generated for other languages but may require additional steps.
-For example, to debug a Rust target program, the Cargo.toml file must list ChatDBG as a dependency and the main function must be annotated with #[chatdbg::main] to ensure that error messages are visible to ChatDBG through a log file.
-IV-C ChatDBG Architecture Overview
-ChatDBG orchestrates communication between the user, the debugger, as shown in the architecture diagram and ChatDBG’s top-level command processing loop in Figure 4.
-The operations in the pseudocode map naturally onto debugger APIs and onto LLM APIs supporting completion and function calls [4].
-ChatDBG currently utilizes OpenAI’s API [17] and GPT-4 models.
-We elaborate on the most salient technical innovations after this overview.
-[…]
-[…]
-Command Debugger Output
-info symbol Pdb The source code and/or docstring for a symbol referring to any function, method, field, class, or package.
-code loc LLDB The source code surrounding loc, where loc is filename:lineno.
-definition loc symbol LLDB The declaration for the first occurrence of symbol at loc, where loc is filename:lineno.
-slice symbol Pdb The source code in the backwards slice of the global symbol. Interactive IPython/Notebook sessions only.
-TABLE II: ChatDBG command extensions. These commands are available to not only the user but also the LLM, and they provide access to information beyond what is typical in a debugger.
-IV-D Initial Prompts and Enriched Stack Traces
+Figure 5. The initial prompt for the debugging session in Figure 2. For brevity, the enriched stack includes only five lines of source in each frame, rather than the default of 10.
+4.4. Initial Prompts and Enriched Stack Traces
 In addition to including the user’s text, the initial prompt conveys instructions to LLM and also the context surrounding the error.
-We highlight the components comprising that context in this section, using the initial prompt in Figure 5 that was generated for the first query in Figure 2.
+We illustrate the components of the prompt in this section, using the initial prompt in Figure 5 that was generated for the first query in Figure 2 as a running example.
 Instructions.
 The instructions at the top of the prompt ask the LLM to answer questions about the root cause of the error, to focus on user code, to explain values stored in variables, and to end each response with either a fix or suggestions for further debugging steps.
-The last item ensures a relatively consistent structure on answers that facilitates reading them and evaluating their quality.
-Paragraphs 2-4 of the instructions are the take the wheel prompt described in Section IV-E.
+The last item ensures a relatively consistent structure for answers that facilitates reading them and evaluating their quality.
+Paragraphs 2-4 of the instructions are the take the wheel prompt described in Section 4.5.
 Enriched stack trace.
@@
 This recursive conversion of values to strings is limited to a depth of three, at which point any remaining values are again abbreviated with ellipses.
+This strategy balances the need to provide the LLM with sufficient information to diagnose the error with the need to avoid overwhelming it with too much information.
+In cases with the elided details are important, the LLM can request them via the take the wheel mechanism.
 ChatDBG follows roughly the same approach in LLDB, utilizing the static types embedded in the DWARF debugging information to decode the stack.
@@
 History. The initial prompt also includes the history of commands already issued by the user, as well as their outputs. This builds a more complete context surrounding the user’s query.
+4.5. Taking the Wheel
+ChatDBG supports take the wheel debugging via the function call capabilities in OpenAI’s API and most recent models (OpenAI, 2023a).
+This feature lets clients register callback functions with the LLM for obtaining additional information while constructing a response.
+The LLM calls these functions by sending special messages to the client as part of its response stream.
+The client receives those messages, computes the requested results, and sends them back to the LLM.
+The initial prompt describes how to use the available functions.
+For example, ChatDBG registers a debug(command) function for running a command in the underlying debugger.
+The LLM calls debug("p len(stats)") through this mechanism in the session from Figure 2.
+ChatDBG then runs Pdb’s command processing routine, onecmd("p len(stats)"), and captures the output to and send back.
+ChatDBG similarly uses the SBCommandInterpreter.HandleCommand routine in LLVM.
+In both cases, the command and output are printed so the user can see these steps.
+The LLM has sufficient background knowledge on debuggers and requires no additional training to navigate up/down the stack, inspect variables and heap data, evaluate expressions, and perform other typical debugger operations.
+Supporting agentic reasoning over run-time program state via function calls is a key technical innovation of ChatDBG.
+Without this capability, there would be no effective way to provide the LLM with a detailed view of relevant program state.
+A common alternative technique for handling large amounts of task-specific data in LLMs is to employ a retrieval augmented generation (RAG) model (DBLP:conf/nips/LewisPPPKGKLYR020), which collects and stores the data in a vectorized database that is then made available to the model for retrieval.
+However, that approach seems less useful in this context, as program state information will be distinct for each debugging session and not easily vectorized.
+Command Debugger Output
+info symbol Pdb The source code and/or docstring for a symbol referring to any function, method, field, class, or package.
+slice symbol Pdb The source code in the backwards slice of the global symbol. Interactive IPython/Notebook sessions only.
+code loc LLDB The source code surrounding loc, where loc has the form filename:lineno.
+definition loc symbol LLDB The declaration for the first occurrence of symbol at loc, where loc has the form filename:lineno.
+Table 2.
+ChatDBG command extensions.
+ChatDBG extends the underlying debuggers with several new commands to help the LLM navigate through and understand the target’s code.
+ChatDBG provides access to them via the LLM’s function call API.
+4.6. Navigating the Code
+While the LLM can often leverage pre-existing background knowledge of common Python and C/C++ standard libraries, it will likely have limited-to-no knowledge of any user-defined code or third-party library functions.
+Trying to include all possibly-relevant source code in the initial prompt would be infeasible and would prevent ChatDBG from scaling to larger codebases.
+Instead, ChatDBG extends the underlying debuggers with several new commands that are designed to help the LLM navigate through and understand the target’s code.
+These commands are available to the LLM via function calls and are listed in Table 2.
+ChatDBG augments Pdb with the info command, which prints the docstring for any function, class, field, method, or package.
+It additionally prints the source for any user-defined code.
+The info requests in Figure 2 demonstrate these two cases for proportion_blue and bootstrap_statistic, respectively.
+The command is implemented via the standard inspect and pydoc libraries.
+The info command is not directly reproducible for unmanaged code in LLVM because there is no comparable existing debugger support for retrieving the source or documentation for a symbol.
+Instead, ChatDBG adds two other debugging commands to LLDB.
+The first, code, prints the code surrounding a source location described by a filename and line number, as in code polymorph.c:118.
+The second command, definition, prints the location and source code for the definition corresponding to the first occurrence of a symbol on a given line of code.
+For example, definition polymorph.c:118 target prints the location and source for the declaration of target corresponding to its use on that line.
+The definition implementation leverages the clangd language server, which supports source code queries via JSON-RPC and Microsoft’s Language Server Protocol (Microsoft, 2016).
+4.7. Slices for Interactive Python
+ChatDBG supports debugging interactive IPython sessions and Jupyter notebooks.
+Interactive sessions lead to many individual code cells that are each evaluated separately.
+Cells may be evaluated out-of-order, override definitions from earlier cells, and communicate values to other cells through top-level global variables.
+Others have noted the challenges of reasoning about program behavior in this context (Head et al., 2019; Shankar et al., 2022).
+ChatDBG provides an additional slice debugging command to facilitate that reasoning.
+The slice command computes the backwards slice for any variable used in the current cell that was defined in previously-executed cells.
+It returns the code for cells in that slice.
+Suppose the code from bootstrap.py in Figure 1 were written in four notebook cells:
+[…]
+[fontsize=]pythonslice.py
+After evaluating these cells, slice(observed_samples) returns the source for the cells labeled In[2] and In[5], and slice(stats) returns the source for all four cells.
+ChatDBG uses ipyflow to compute slices (Shankar et al., 2022; IPyflow, 2020).
+4.8. Security and Risks
+It is possible for the LLM to issue debugging commands containing arbitrary code through the debug function call provided by ChatDBG.
+That code could, for example, delete files or execute other malicious actions on the client.
+ChatDBG mitigates this risk by sanitizing LLM-generated debugging commands before running them.
+For Python, the sanitizer ensures that any functions called in LLM-provided commands belong to a user-configurable whitelist.
+For native code, code provenance is harder to track and languages are more permissive, so the sanitizer rejects any commands calling functions.
+ChatDBG supports an --unsafe flag to disable sanitizing when the client system is running in an isolated environment that obviates the need for such protections.
+It is also possible for the LLM to hallucinate and respond with incorrect or misleading diagnoses and fixes.
+ChatDBG mitigates this risk by not directly applying proposed code fixes or suggestions to the target code.
+Instead, ChatDBG presents them to the user, who may then vet and judge the quality of the LLM’s responses and decide whether or not to follow suggested changes.
 Name LoC Type Reported Exception Root Cause
@@
 s10 118 semantic Assertion Error Loops doesn’t append to array correctly
-s11 181 crash Value Error Takes a sample from a table smaller than the sample size without replacement
+s11 181 crash Value Error Creates a sample without replacement that is larger than the input
 s12 127 crash Value Error Incorrect label when accessing column value for table row
 s13 127 crash Value Error Pivot uses wrong columns for row/columns in new table
-s14 65 crash Index Error Simulation fails to properly create a random sample under null hypothesis
-TABLE III: Python programs exhibiting a variety of common errors.
+s14 65 crash Index Error Incorrect computation of a random sample under null hypothesis
+Table 3.
+Python programs exhibiting a variety of common errors.
 Programs c1–c8 are command line scripts, and programs s1–s14 are Jupyter notebooks, which utilize two non-standard libraries consisting of 3,000 lines of code.
@@
 Crashes reflect unexpected termination due to any other type of error.
-IV-E Taking the Wheel
-ChatDBG supports take the wheel debugging via the function call capabilities in OpenAI’s API and most recent models [4].
-This feature lets clients register callback functions with the LLM for obtaining additional information while constructing a response.
-The LLM calls these functions by sending special messages to the client as part of its response stream.
-The client receives those messages, computes the requested results, and sends them back to the LLM.
-The initial prompt describes how to use the available functions.
-For example, ChatDBG registers a debug(command) function for running a command in the underlying debugger.
-The LLM calls debug("p len(stats)") through this mechanism in the session from Figure 2.
-ChatDBG then runs Pdb’s command processing routine, onecmd("p len(stats)"), and captures the output to and send back.
-ChatDBG similarly uses the SBCommandInterpreter.HandleCommand routine in LLVM.
-In both cases, the command and output are printed so the user can see these steps.
-The LLM has sufficient background knowledge on debuggers and requires no additional training to navigate up/down the stack, inspect variables and heap data, evaluate expressions, and perform other typical debugger operations.
-IV-F Navigating the Code
-While the LLM can often leverage pre-existing background knowledge of common Python and C/C++ standard libraries, it will likely have limited-to-no knowledge of any user-defined code or third-party library functions.
-ChatDBG extends the underlying debuggers with several new commands that are designed to help the LLM navigate through and understand the target’s code.
-These commands are available to the LLM via function calls and are listed in Figure II.
-ChatDBG augments Pdb with the info command, which prints the docstring for any function, class, field, method, or package.
-It additionally prints the source for any user-defined code.
-The info requests in Figure 2 demonstrate these two cases for proportion_blue and bootstrap_statistic, respectively.
-The command is implemented via the standard inspect and pydoc libraries.
-The info command is not directly reproducible for unmanaged code in LLVM because there is no comparable existing debugger support for retrieving the source or documentation for a symbol.
-Instead, ChatDBG adds two other debugging commands to LLDB.
-The first, code, prints the code surrounding a source location described by a filename and line number, as in code polymorph.c:118.
-The second command, definition, prints the location and source code for the definition corresponding to the first occurrence of a symbol on a given line of code.
-For example, definition polymorph.c:118 target prints the location and source for the declaration of target corresponding to its use on that line.
-The definition implementation leverages the clangd language server, which supports source code queries via JSON-RPC and Microsoft’s Language Server Protocol [18].
-IV-G Slices for Interactive Python
-ChatDBG supports debugging interactive IPython sessions and Jupyter notebooks.
-Interactive sessions lead to many individual code cells that are each evaluated separately.
-Cells may be evaluated out-of-order, override definitions from earlier cells, and communicate values to other cells through top-level global variables.
-Others have noted the challenges of reasoning about program behavior in this context [19, 20].
-ChatDBG provides an additional slice debugging command to facilitate that reasoning.
-The slice command computes the backwards slice for any variable used in the current cell that was defined in previously-executed cells.
-It returns the code for cells in that slice.
-Suppose the code from bootstrap.py in Figure 1 were written in four notebook cells:
-In[2]: def make_marble_sample(): ...
-In[3]: def proportion_blue(sample): ...
-In[4]: def resampled_stats(observed_marbles, num_trials):
-           stats = bootstrap_statistic(observed_marbles,
-                                       proportion_blue,
-                                       num_trials)
-           assert len(stats) == num_trials
-           return stats
-In[5]: observed_marbles = make_marble_sample()
-       stats = resampled_stats(observed_marbles, 5)
-After evaluating these cells, slice(observed_samples) returns the source for the cells labeled In[2] and In[5], and slice(stats) returns the source for all four cells.
-ChatDBG uses ipyflow to compute slices [20, 21].
-V Evaluation
+5. Evaluation
 We demonstrate ChatDBG’s capacity to identify the root cause of defects and provide fixes in two contexts: bugs in relatively small Python programs written by students and bugs in large C/C++ programs.
@@
 RQ3: Is ChatDBG effective at diagnosing and fixing bugs in unmanaged code (C/C++)?
-V-A Python
-We applied ChatDBG to bugs in a collection of student labs from two introductory computer science courses; see Table III.
+5.1. Python
+We applied ChatDBG to bugs in a collection of student labs from two introductory computer science courses; see Table 3.
 Bugs c1–c8 are in non-interactive scripts from a programming class that perform various file reading and text processing tasks.
-Bugs s1–s14 are in Jupyter notebooks [22] from a data science class that manipulate, visualize, and compute over arrays and tables.
+Bugs s1–s14 are in Jupyter notebooks (Team, 2015) from a data science class that manipulate, visualize, and compute over arrays and tables.
 Some bugs were apparent to the programs’ authors.
@@
 Programs were prepared by removing them from their automatic grading harness and replacing failed unit tests with assert statements that generate exceptions.
-We ran each program ten times under the five configurations in Table IV: Default Stack includes standard stack traces, as generated by ipdb [23], with 5 lines of code per frame in the initial prompt, but it does not support the LLM taking the wheel.
+We ran each program ten times under the five configurations in Table 4: Default Stack includes standard stack traces, as generated by ipdb (ipdb, 2007), with 5 lines of code per frame in the initial prompt, but it does not support the LLM taking the wheel.
 Enriched Stack generates enriched stacks with ten lines of code per frame, and +Take the Wheel additionally permits ChatDBG to run debugger commands.
@@
 For crashes, the questions relate the crash to expected behavior, as in the following; we designed our questions to be “neutral” and not hint at the root cause.
-c3 (Crash)
+c3 (Crash):
 Why am I not reading the CSV file correctly?
-s11 (Crash)
+s11 (Crash):
 Why am I not able to sample 100 rows?
-c1 (Semantic)
+c1 (Semantic):
 Why am I not getting 3?
-s1 (Semantic)
+s1 (Semantic):
 bill_length_mean_by_species should be a table of the mean bill lengths of each species in our data set. Why isn’t it?
@@
 ChatDBG used the gpt-4-1106-preview model for these experiments.
-Under +Targeted Question, the first prompt and response led to, on average, a chat of about 10,000 tokens (7,500 words), a cost of about \$$0.120.12$ USD under OpenAI’s current pricing model [24], and a completion time of about 25 seconds.
+Under +Targeted Question, the first prompt and response led to, on average, a chat of about 10,000 tokens (7,500 words), a cost of about \$$0.120.12$ USD under OpenAI’s current pricing model (OpenAI, [n.
+d.]), and a completion time of about 25 seconds.
 Subsequent steps in extended debugging dialogs incurred comparable costs.
 Time was highly variable and dominated by the performance of OpenAI’s service.
-These characteristics will be different for other platforms and models.
-RQ1: Is ChatDBG effective at diagnosing and fixing bugs in Python?
-Each response was manually examined and deemed a success if it included an accurate explanation of the error and an actionable fix.
-That fix could be either code or a prose description in which all necessary details were made explicit.
-Figure 6 shows the success rate under each configuration.
-RQ1 Summary: Even with just the simple question why?, ChatDBG was successful 57% of the time.
-With questions specialized to the target’s particular error, that number jumps to 67%, and with an additional dialog step ChatDBG succeeded in identiyfing and fixing the defect in 85% of the trials.
+These characteristics will be different for other platforms and models and, given current trends, we expect significant reductions in both time and cost as models improve.
 Configuration Stack Take the Initial Ask a
@@
 +Dialog enriched specialized
-TABLE IV: Configurations used in the Python experiments.
-Figure 6: Overall ChatDBG success rate for each configuration. ChatDBG innovations and user-provided context gradually increase effectiveness.
-Figure 7: Success rate for ChatDBG for each program and configuration. Vertical lines show the mean.
-Program LoC Error Type Root Cause   Proximate Cause Fix Root Cause Fix
-BC [25] 17.0k Buffer overflow Input from data file printed to a fixed-size buffer Truncate on copy Use dynamic size
-GZIP [25] 8.2k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
-NCOM [25] 1.9k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
-PEG [26] 14.7k Null dereference Invalid input produces corrupted data structure Check if not null Warn/exit
-POLY [25] 0.7k Buffer overflow Command line argument is unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
-TIFF [26] 58.9k Division by zero Combination of command line options leads to a division by zero Override option values Warn/exit when invalid
-YAML1 [26] 8.7k Stack overflow Long sequences of { in the input leads to deep recursion Use iterative method Guard recursion depth
-YAML2 [26] 8.7k Assertion failure Specific input causes a peek request for non-existent “next” token Replace assert Check before peeking
-TABLE V: Bugs in unmanaged C/C++ code, and our criteria for fixing the proximate cause or the root cause of each.
+Table 4. Configurations used in the Python experiments.
+Figure 6. Overall ChatDBG success rate for each configuration. ChatDBG innovations and user-provided context gradually increase effectiveness.
+RQ1: Is ChatDBG effective at diagnosing and fixing bugs in Python?
+Each response was manually examined and deemed a success if it included an accurate explanation of the error and an actionable fix.
+That fix could be either code or a prose description in which all necessary details were made explicit.
+To avoid bias in this assessment, explicit criteria for each program was determined prior to examining the responses.
+Figure 6 shows the success rate under each configuration.
+With all features enabled, ChatDBG was successful at identifying and fixing bugs in well over half of the trials.
+Any time or energy expended by the user manually debugging those cases would be all but eliminated by using ChatDBG.
+RQ1 Summary: Even with just the simple question why?, ChatDBG was successful 57% of the time.
+With questions specialized to the target’s particular error, that number jumps to 67%, and with an additional dialog step ChatDBG succeeded in identifying and fixing the defect in 85% of the trials.
 RQ2: Which components of ChatDBG contribute to its effectiveness?
 Figure 7 presents the success rates for each program under each configuration.
-The Enriched Stack plots demonstrate that enriched stacks provide some benefit, particuarly for crashes in which the stack contains sufficient information to diagnose the problem, but they alone do not provide much improvement for many semantic errors in which the relevant computation steps complete before failure.
+The Enriched Stack plots demonstrate that enriched stacks provide some benefit, particularly for crashes in which the stack contains sufficient information to diagnose the problem, but they alone do not provide much improvement for many semantic errors in which the relevant computation steps complete before failure.
 However, enriched stacks coupled with letting the LLM taking the wheel led to significant improvement in the success rate for both crashing and semantic bugs, as shown in the +Take the Wheel plots.
@@
 Without it, success rates rarely improved when the LLM took the wheel.
-The +Targeted Question configuration demonstrates the impact of providing even the most modest details about expected behavior.
+Figure 7. Success rate for ChatDBG for each program and configuration. Vertical lines show the mean.
+Error Fix
+Program LoC Type Root Cause Proximate Cause Root Cause
+BC (Lu et al., 2005) 17.0k Buffer overflow Input from data file printed to a fixed-size buffer Truncate on copy Use dynamic size
+GZIP (Lu et al., 2005) 8.2k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
+NCOM (Lu et al., 2005) 1.9k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
+PEG (An et al., 2023) 14.7k Null dereference Invalid input produces corrupted data structure Check if not null Warn/exit
+POLY (Lu et al., 2005) 0.7k Buffer overflow Command line argument is unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
+TIFF (An et al., 2023) 58.9k Division by zero Combination of command line options leads to a division by zero Override option values Warn/exit when invalid
+YAML1 (An et al., 2023) 8.7k Stack overflow Long sequences of { in the input leads to deep recursion Use iterative method Guard recursion depth
+YAML2 (An et al., 2023) 8.7k Assertion failure Specific input causes a peek request for non-existent “next” token Replace assert Check before peeking
+Table 5. Bugs in unmanaged C/C++ code, and our criteria for fixing the proximate cause or the root cause of each.
+The +Targeted Question configuration demonstrates the impact of providing even the most modest details about expected behavior in queries.
 When the LLM is asked to continue its reasoning in +Dialog, ChatDBG’s success rate improves despite the follow-up prompt providing no feedback on the contents or quality of the first response.
-This phenom indicates that constraints on the underlying LLM’s response lengths may prevent it from conducting the amount of reasoning necessary to successfully develop a fix in a single step.
+This phenom indicates that constraints on the underlying LLM’s response lengths may prevent it from conducting the amount of reasoning necessary to develop a fix in a single step.
 The success rates for +Targeted Question and +Dialog demonstrate the importance of continued dialogs and user input.
 We expect those features to be even more important to ChatDBG’s success when diagnosing bugs in more complex programs.
-The LLM also demonstrated its background knowledge with the responses including, for example, details of Python idioms and libraries, the definition of h-index [27], and the implementation and limitations of various statistical techniques.
+The LLM also demonstrated its background knowledge with the responses including, for example, details of Python idioms and libraries, the definition of h-index (Hirsch, 2005), and the implementation and limitations of various statistical techniques.
 Failures were generally due to the LLM not always recognizing or discovering key aspects of a program’s behavior.
+In general, the further the distance between the root cause of a bug and observable effect, the more challenging it was for ChatDBG (and people!) to find it.
 In some cases, it was on the right track but did not converge on an actionable fix.
@@
 The most sophisticated configurations show that user-provided contextual information about behavior and engaging in multi-step dialogs are particularly good ways to improve its effectiveness.
-V-B C and C++
+5.2. C and C++
 Programs in unmanaged languages such as C and C++ are vulnerable to memory safety errors.
 These memory errors can also hinder the debugging process: the crash may not occur immediately at the memory violation but instead much later on, and the crash may cause corruption of the stack and/or heap, making it challenging to recover any useful information.
-Table V summarizes the programs from the BugBench [25] and BugsC++ [26] suites used to evaluate ChatDBG’s effectiveness at debugging unmanaged code.
+Table 5 summarizes the programs extracted from the BugBench (Lu et al., 2005) and BugsC++ (An et al., 2023) suites used to evaluate ChatDBG’s effectiveness at debugging unmanaged code.
 Programs used in this evaluation are all real-world applications with concrete known bugs.
@@
 Some of the studied programs do not crash at run time.
-We employed AddressSanitizer [28] to force a crash at the moment a memory violations occured to trigger those defects.
+We employed AddressSanitizer (Serebryany et al., 2012) to force a crash at the moment a memory violation occurs to trigger those defects.
 AddressSanitizer is already capable of reporting some information about the crash when it happens.
@@
 RQ3: Is ChatDBG effective at diagnosing and fixing bugs in unmanaged code (C/C++)?
-We run our C/C++ experiments on an an x86 Ubuntu 22.04 server.
-We use Clang and LLDB 17.0.6 to compile and debug, using flags -g -Og -fno-omit-frame-pointer.
+We ran our C/C++ experiments on an an x86 Ubuntu 22.04 server.
+We used Clang and LLDB 17 to compile and debug, using flags -g -Og -fno-omit-frame-pointer.
 ChatDBG used OpenAI’s gpt-4-1106-preview model.
@@
 We manually examined each response to determine if ChatDBG successfully provided an actionable code fix for the proximate cause of the crash or for the underlying root cause.
-We used the critera outlined in Table V.
+We used the criteria outlined in Table 5.
 While fixing root causes is the ultimate goal, fixing proximate causes can still be beneficial as fixing crashes enables further debugging steps.
-Figure 8: ChatDBG success rate at fixing the proximate or root cause in C/C++ programs.
+Figure 8.
+ChatDBG success rate at fixing the proximate or root cause in C/C++ programs.
 ChatDBG successfully identified and fixed the root cause 36% of the time and the proximate cause an additional 55% of the time.
@@
 It was also capable of providing relevant, actionable fixes: 36% of its suggestions addressed the root cause of the bug, while another 55% corrected the proximate cause.
-V-C Threats to Validity
+5.3. Threats to Validity
 This paper evaluates ChatDBG on two suites of code.
@@
 While the C/C++ suite consists of real-world applications, most of the errors are memory errors.
-Other types of errors, such as assertion failures, concurrency errors, or other logical errors, may lead to different results.
+Other types, such as assertion failures, concurrency errors, or other logical errors, may lead to different results.
 ChatDBG depends on an LLM to analyze and drive exploration of state, and like all systems based on LLMs today, its performance is affected by prompt engineering.
@@
 To mitigate this threat, our evaluation runs ChatDBG on each test program at least ten times.
-Our evaluation depends on a manual and subjective evaluation of whether ChatDBG’s explanation of a bug and its proposed fix are satisfactory.
-We mitigated the risks of subjective evaluation by using precisely-defined criteria decided upon in advance.
+Our evaluation depends on a manual evaluation of whether ChatDBG’s explanation of a bug and its proposed fix are satisfactory.
+We mitigated the risks of subjectivity by using precisely-defined criteria decided upon in advance.
 Python fixes were deemed successful if the resulting code met the correctness requirements outlined in the assignment.
-C/C++ fixes were deemed successful at fixing the proximate or root cause using the criteria in Table V.
+C/C++ fixes were deemed successful at fixing the proximate or root cause using the criteria in Table 5.
 Fixes described in prose were permitted, provided that the details of all necessary changes to the code were made explicit.
-VI Conclusion and Future Work
+6. Future Work
+We see several promising avenues for future work.
+Incorporating existing fault localization approaches into ChatDBG, rather than relying solely on the LLM’s ability to explore the program’s code and state, could potentially increase its effectiveness and efficiency by allowing the LLM to focus its attention on suspicious files, functions, or lines of source code.
+Similarly, incorporating delta debugging (Zeller, 1999) could increase the effectiveness of ChatDBG by limiting the amount of input for an LLM and providing failure-inducing events as guidance.
+Finally, integrating ChatDBG with a time-travel debugger (Project, 2009; O’Callahan et al., 2017) would expand its reach to exploring program state over time, letting it answer queries that cannot be answered given only the current program state.
+One challenge of integrating these more sophisticated techniques will be ensuring that the LLM can effectively utilize them, which may necessitate fine tuning or additional training on their usage.
+7. Conclusion
 This paper presents ChatDBG, the first AI-based debugging assistant.
 Our evaluation shows that engaging in a debugging dialog with ChatDBG can significantly assist in identifying root causes of errors and developing correct fixes.
-We see several avenues for future work.
-Incorporating an existing fault localization approach into ChatDBG, rather than relying solely on the LLM’s ability to explore the program’s source code and state, could potentially increase its effectiveness and efficiency by allowing the LLM to focus its attention on suspicious and possibly problematic files, functions, or lines of source code.
-Similarly, incorporating delta debugging [29] could increase the effectiveness of ChatDBG by limiting the amount of input for an LLM and providing failure-inducing events as guidance.
-Finally, integrating ChatDBG with a time-travel debugger would expand its reach to exploring program state over time, letting it answer queries that cannot be answered given the current program state.
-ChatDBG is available on GitHub at github.com/plasma-umass/ChatDBG.

### v3

@@
-ChatDBG: An AI-Powered Debugging Assistant
+ChatDBG: Augmenting Debugging with Large Language Models
 Abstract.
@@
 It then reports its findings and yields back control to the programmer.
+By leveraging the real-world knowledge embedded in LLMs, ChatDBG can diagnose issues identifiable only through the use of domain-specific reasoning.
 Our ChatDBG prototype integrates with standard debuggers including LLDB and GDB for native code and Pdb for Python.
@@
 For the Python programs, a single query led to an actionable bug fix 67% of the time; one additional follow-up query increased the success rate to 85%.
-ChatDBG has seen rapid uptake; it has already been downloaded roughly 50,000 times.
+ChatDBG has seen rapid uptake; it has already been downloaded more than 65,000 times.
 Debugging, Artificial Intelligence, Software Engineering
-$c$s: Computing methodologies Artificial intelligenceccs: Software and its engineering Software testing and debuggingconference: Under Submission; 2025; Someplace, Somewhere
+$c$s: Computing methodologies Artificial intelligenceccs: Software and its engineering Software testing and debugging
 1. Introduction
@@
 data breakpoints); inspecting local variables, globals, heap objects, and backtraces of the call stack; and resuming program execution line-by-line (single-step) or at the granularity of function calls.
-Debuggers can be helpful, but finding and fixing software defects remains a deeply challenging and time-consuming task.
+Debuggers can be helpful, but finding and fixing software defects remains a deeply challenging and time-consuming task (Zeller, 2009a; Eisenstadt, 1993; Layman et al., 2013).
 Programmers must still reason about program behavior to ascertain what went wrong.
@@
 Such information includes lengthy executions, large amounts of program data, and many stack frames that potentially span multiple threads.
-This paper introduces ChatDBG, the first AI-powered debugger assistant.
+This paper introduces the ChatDBG AI-powered debugger assistant.
 ChatDBG integrates into and significantly extends the functionality of standard debuggers.
 ChatDBG builds on the insight that large language models (LLMs), such as OpenAI’s GPT-4 (OpenAI, 2023b), enable a debugger to leverage insights and intuition from the vast real-world knowledge embedded in LLMs.
+This knowledge enables ChatDBG to fix classes of issues that depend on logical thinking and domain-specific reasoning beyond the ability to write and debug programs.
+For example, Figure 2 illustrates the use of ChatDBG to debug a program by leveraging a knowledge of statistics that cannot be gleaned from the program itself.
 A debugger integrated with ChatDBG continues to provide its full range of functionality but also lets programmers engage in debugging dialogs where they can ask high-level questions like ‘why is x null here?’ or ‘why isn’t this value what I expected?’.
@@
 To answer such queries, ChatDBG orchestrates a conversation with an LLM.
-A key novelty of ChatDBG is that it grants autonomy to the LLM to “take the wheel” and act as an independent agent while answering the programmer’s queries.
+A key feature of ChatDBG is that it grants autonomy to the LLM to “take the wheel” and act as an independent agent (Wooldridge and Jennings, 1995; Gur et al., 2024) while answering the programmer’s queries.
 Specifically, the LLM issues “function calls” (OpenAI, 2023a) to run commands in the underlying debugger to investigate program state, execute code, or obtain source code.
@@
 After answering a query, control is returned to the programmer, who may then enter additional commands or chat messages.
-Support for the LLM’s agentic control of the debugger is a key technical contribution.
 Our prototype of ChatDBG integrates into three widely used debuggers: GDB, LLDB, and Pdb.
@@
 This paper makes the following contributions.
-  • It introduces ChatDBG, the first AI-powered debugger assistant.
+  • It introduces ChatDBG, an AI-powered debugger assistant that enables large language models to “take the wheel” and control the debugger via agentic reasoning.
   • It describes the implementation of our ChatDBG prototype.
-  • It introduces the “take the wheel” approach to integrating agentic reasoning in large language models with developer tools.
   • It presents an evaluation of ChatDBG that demonstrates its significant advantages over existing debugger functionality.
-ChatDBG is broadly applicable to many domains and programming languages, and we expect it to be particularly useful for novice programmers, who often lack the experience to effectively use debuggers.
-ChatDBG is also useful for experienced programmers, who can augment debugging sessions with ChatDBG’s reasoning capabilities in a natural, conversational, and interactive way.
+Our evaluation shows that ChatDBG is broadly applicable to many domains and programming languages, and we expect it to be particularly useful for novice programmers, who often lack the experience to effectively use debuggers.
+ChatDBG is also useful for experienced programmers, who can augment debugging sessions with ChatDBG’s reasoning capabilities in a conversational and interactive way.
 This paper is organized as follows.
@@
 Figure 1.
-An example program containing several bugs.
+An example program containing several bugs (§2).
 It is supposed to create an array of marble colors, compute the proportions of blue marbles in resamples of that array, and assert that their mean is about 0.7, the proportion for the array.
@@
 Figure 2.
-A debugging session with ChatDBG.
+A debugging session with ChatDBG2).
 For brevity, we elide repetitive or unimportant parts.
@@
 Figure 3.
-A debugging session demonstrating how ChatDBG incorporates real-world knowledge.
+A debugging session demonstrating how ChatDBG incorporates real-world knowledge (§2).
 After correcting proportion_blue in Figure 1, the program fails on line 21 because the mean proportion of blue marbles in the resamples is not the expected value.
@@
 Table 1.
-Debugger features and their dates of introduction.
+Debugger features and their dates of introduction (§3).
 Most key features have been around for decades.
@@
 DeepFL (Li et al., 2019), Grace (Lou et al., 2021) and DeepRL4FL (Li et al., 2021) are examples of such systems.
-Similarly, LLMAO (Yang et al., 2024) employs LLMs to provide suspiciousness scores for each line of code in a given program, but only provides access to the source code.
+Similarly, LLMAO (Yang et al., 2024b) employs LLMs to provide suspiciousness scores for each line of code in a given program, but only provides access to the source code.
 AutoFL (Kang et al., 2024) also utilizes an LLM and enables it to statically retrieve source code and coverage information about the program via function calls.
@@
 ChatDBG improves upon these systems by providing an LLM with access to run-time program state and the ability to take control of the underlying debugger.
-Both of these features enhance the LLM’s ability to provide more accurate and informative feedback to the user.
+Both features enhance the LLM’s ability to provide more accurate and informative feedback to the user.
 We also note that other fault localization techniques can be used in tandem with ChatDBG to improve results, as suggested by Section 5.1’s utilization of backwards slicing to identify code relevant to a bug in Python notebooks and Section 5.2’s utilization of AddressSanitizer (Serebryany et al., 2012) to provide a better starting point for diagnosing and fixing memory errors in native code.
@@
 Previous research has shown that automated program repair hints can provide significant help in the debugging process and suggests that the benefits of correct advice outweigh the risk of deceptive ones (Eladawy et al., 2024).
+3.1. Concurrent work
+Several approaches developed concurrently with ChatDBG have also integrated LLMs into automatic program repair or fault localization techniques to enhance the debugging process.
+Robin (Bajpai et al., 2024) is a chat-based debugging assistant designed to help users diagnose errors more quickly.
+Both Robin and ChatDBG provide a limited program context to the LLM at the beginning of a conversation.
+However, Robin has no direct access to any additional context about the program and execution state; the user must manually retrieve and provide these items.
+Robin’s functionality is therefore roughly equivalent to the Enriched Stack configuration of ChatDBG, detailed in Section 5.1.
+As Figure 6 shows, ChatDBG achieves a nearly two-fold increase improvement in diagnosing errors versus the Enriched Stack configuration.
+ChatDBG’s effectiveness generally increases further with targeted questions and follow-up discussions with the user.
+AutoCodeRover (Zhang et al., 2024) and SWE-agent (Yang et al., 2024a) are complementary approaches that focus on fault localization and automatic repair, relying exclusively on issue descriptions and source code.
+ChatDBG additionally leverages run-time information to identify root causes and propose fixes.
+Section 5 demonstrates the strength of this approach over relying solely on static information.
+Both AutoCodeRover and SWE-agent perform an evaluation using SWE-bench (Jimenez et al., 2024), which was created to evaluate the efficacy of such static tools; unfortunately, this benchmark suite is not applicable to ChatDBG due to its extensive usage of run-time information.
+RepairAgent (Bouzenia et al., 2024) and AutoSD (Kang et al., 2025) are tools that employ LLMs in specific workflows that mimic standard debugging strategies in an attempt to repair pre-identified bugs.
+While successful in some settings, both tools rely on the user providing a failing test case and the precise location of the bug.
+By contrast, ChatDBG does not require this information.
+ChatDBG also enables a more flexible workflow that permits collaboration with the developer in addition to seamless integration into the standard debugging process.
 4. Implementation
@@
 […]
-Figure 4. ChatDBG architecture and command processing algorithm.
+Figure 4. ChatDBG architecture and command processing algorithm (§4.3).
 4.3. ChatDBG Architecture Overview
@@
 Why doesn’t stats have 5 elements?
-Figure 5. The initial prompt for the debugging session in Figure 2. For brevity, the enriched stack includes only five lines of source in each frame, rather than the default of 10.
+Figure 5. The initial prompt for the debugging session in Figure 24.4). For brevity, the enriched stack includes only five lines of source in each frame, rather than the default of 10.
 4.4. Initial Prompts and Enriched Stack Traces
-In addition to including the user’s text, the initial prompt conveys instructions to LLM and also the context surrounding the error.
+In addition to including the user’s text, the initial prompt conveys instructions to LLM and the context surrounding the error.
 We illustrate the components of the prompt in this section, using the initial prompt in Figure 5 that was generated for the first query in Figure 2 as a running example.
@@
 Similarly, ChatDBG recursively converts the values stored in aggregate structures like lists, arrays, and dictionaries to strings, but limits the number of elements shown to a small, fixed number.
-The rest of the elements are just abbreviated with an ellipsis (…).
+The rest of the elements are abbreviated with an ellipsis (…).
 This recursive conversion of values to strings is limited to a depth of three, at which point any remaining values are again abbreviated with ellipses.
@@
 Error.
-A description of the error causing execution to stop is also extracted from the underlying debugger.
+A description of the error causing execution to stop is extracted from the underlying debugger.
 When the error is due to an assertion failure, ChatDBG instructs the LLM to assume that the assertion is valid as written so that it will look beyond the assertion for the real problem.
@@
 ChatDBG supports take the wheel debugging via the function call capabilities in OpenAI’s API and most recent models (OpenAI, 2023a).
-This feature lets clients register callback functions with the LLM for obtaining additional information while constructing a response.
+This agentic approach (Wooldridge and Jennings, 1995; Gur et al., 2024) lets clients register callback functions with the LLM for obtaining additional information while constructing a response.
 The LLM calls these functions by sending special messages to the client as part of its response stream.
@@
 Without this capability, there would be no effective way to provide the LLM with a detailed view of relevant program state.
-A common alternative technique for handling large amounts of task-specific data in LLMs is to employ a retrieval augmented generation (RAG) model (DBLP:conf/nips/LewisPPPKGKLYR020), which collects and stores the data in a vectorized database that is then made available to the model for retrieval.
+A common alternative technique for handling large amounts of task-specific data in LLMs is to employ a retrieval augmented generation (RAG) model (Lewis et al., 2020), which collects and stores the data in a vectorized database that is then made available to the model for retrieval.
 However, that approach seems less useful in this context, as program state information will be distinct for each debugging session and not easily vectorized.
@@
 Table 2.
-ChatDBG command extensions.
+ChatDBG command extensions (§4.6).
 ChatDBG extends the underlying debuggers with several new commands to help the LLM navigate through and understand the target’s code.
@@
 Trying to include all possibly-relevant source code in the initial prompt would be infeasible and would prevent ChatDBG from scaling to larger codebases.
-Instead, ChatDBG extends the underlying debuggers with several new commands that are designed to help the LLM navigate through and understand the target’s code.
-These commands are available to the LLM via function calls and are listed in Table 2.
+Instead, ChatDBG extends the underlying debuggers with several new commands designed to help the LLM navigate through and understand the target’s code.
+These commands are available to the LLM via function calls and listed in Table 2.
 ChatDBG augments Pdb with the info command, which prints the docstring for any function, class, field, method, or package.
@@
 Table 3.
-Python programs exhibiting a variety of common errors.
+Python programs exhibiting a variety of common errors (§5.1).
 Programs c1–c8 are command line scripts, and programs s1–s14 are Jupyter notebooks, which utilize two non-standard libraries consisting of 3,000 lines of code.
@@
 We demonstrate ChatDBG’s capacity to identify the root cause of defects and provide fixes in two contexts: bugs in relatively small Python programs written by students and bugs in large C/C++ programs.
-The former have well-defined expected behavior that enable us to thoroughly and systematically assess ChatDBG.
+The former have well-defined expected behavior that enables us to thoroughly and systematically assess ChatDBG.
 The latter demonstrates its effectiveness on unmanaged code when unusual corner cases trigger crashes.
@@
 5.1. Python
-We applied ChatDBG to bugs in a collection of student labs from two introductory computer science courses; see Table 3.
+We applied ChatDBG to all bugs in a collection of student labs from two introductory computer science courses; see Table 3.
 Bugs c1–c8 are in non-interactive scripts from a programming class that perform various file reading and text processing tasks.
@@
 In addition, the programs have clear correctness criteria that lead to objective effectiveness metrics in our experiments.
-The bugs are also representative of mistakes often made in languages like Python.
+The bugs are representative of common mistakes because they were introduced by real humans, rather than synthetically generated.
 They range from scoping issues, algorithmic errors, and misuse of library functions to subtle misunderstanding of domain knowledge.
@@
 Programs were prepared by removing them from their automatic grading harness and replacing failed unit tests with assert statements that generate exceptions.
+We focus our evaluation on Python and perform an ablation study by progressively enabling ChatDBG features.
 We ran each program ten times under the five configurations in Table 4: Default Stack includes standard stack traces, as generated by ipdb (ipdb, 2007), with 5 lines of code per frame in the initial prompt, but it does not support the LLM taking the wheel.
@@
 +Targeted Question asks a question specific to the failure.
-For semantic errors, which validate the values stored in variables, these questions describe what those values should be or what they are intended to represent.
+For semantic errors, which validate the values stored in variables, these questions describe what those values should be or what they intend to represent.
 For crashes, the questions relate the crash to expected behavior, as in the following; we designed our questions to be “neutral” and not hint at the root cause.
@@
 +Dialog enriched specialized
-Table 4. Configurations used in the Python experiments.
-Figure 6. Overall ChatDBG success rate for each configuration. ChatDBG innovations and user-provided context gradually increase effectiveness.
+Table 4. Configurations used in the Python experiments (§5.1).
+Figure 6. Overall ChatDBG success rate for each configuration (§5.1). ChatDBG innovations and user-provided context gradually increase effectiveness.
 RQ1: Is ChatDBG effective at diagnosing and fixing bugs in Python?
@@
 Figure 6 shows the success rate under each configuration.
+The simplest configuration, Default Stack, provides functionality roughly equivalent to the user copying and pasting the program stack trace and basic error information into an LLM chat window and requesting a fix.
+We use this configuration as a baseline for evaluating the impact of ChatDBG’s more advanced configurations.
 With all features enabled, ChatDBG was successful at identifying and fixing bugs in well over half of the trials.
@@
 The Enriched Stack plots demonstrate that enriched stacks provide some benefit, particularly for crashes in which the stack contains sufficient information to diagnose the problem, but they alone do not provide much improvement for many semantic errors in which the relevant computation steps complete before failure.
-However, enriched stacks coupled with letting the LLM taking the wheel led to significant improvement in the success rate for both crashing and semantic bugs, as shown in the +Take the Wheel plots.
-The LLM most heavily used the info, slice (for notebooks), and p (print) debugging commands.
-That is not surprising, given that they often provide the most direct insight into the execution state and code.
-The slice command was critically important for notebooks.
+However, enriched stacks coupled with letting the LLM take the wheel led to significant improvement in the success rate for both crashing and semantic bugs, as shown in the +Take the Wheel plots.
+Using the +Take the Wheel feature, the LLM issues from 0 to 12 debugging commands per run, most commonly calling the info, slice (for notebooks), and p (print) debugging commands.
+While all of these commands provide useful information about execution state and code, the slice command was critically important for notebooks.
 Without it, success rates rarely improved when the LLM took the wheel.
-Figure 7. Success rate for ChatDBG for each program and configuration. Vertical lines show the mean.
+Figure 7. Success rate for ChatDBG for each program and configuration (§5.1). Vertical lines show the mean.
 Error Fix
@@
 YAML2 (An et al., 2023) 8.7k Assertion failure Specific input causes a peek request for non-existent “next” token Replace assert Check before peeking
-Table 5. Bugs in unmanaged C/C++ code, and our criteria for fixing the proximate cause or the root cause of each.
+Table 5. Bugs in unmanaged C/C++ code, and our criteria for fixing the proximate cause or the root cause of each (§5.2).
 The +Targeted Question configuration demonstrates the impact of providing even the most modest details about expected behavior in queries.
@@
 Failures were generally due to the LLM not always recognizing or discovering key aspects of a program’s behavior.
-In general, the further the distance between the root cause of a bug and observable effect, the more challenging it was for ChatDBG (and people!) to find it.
+We observe that in some cases, enabling additional features in ChatDBG decreases its success rate.
+We attribute this result to the fact that longer and more complex prompts can occasionally degrade the effectiveness of LLMs (Levy et al., 2024).
+In general, the further the distance between the root cause of a bug and observable effect, the more challenging it was for ChatDBG (and people (Zeller, 2009b, p.
+243)) to find it.
 In some cases, it was on the right track but did not converge on an actionable fix.
@@
 It also occasionally made mistakes, such as conflating proportions and percentages or failing to handle unusual corner cases.
-All of these could be mitigated by feedback from the user in subsequent follow ups.
+All of these could be mitigated by feedback from the user in subsequent follow-ups.
 RQ2 Summary: While all features of ChatDBG contribute to its success, the technical innovations enabling it to take the wheel are critical.
@@
 Programs used in this evaluation are all real-world applications with concrete known bugs.
-The four BugBench programs were selected as the only ones we could retrieve, build and reproduce on our system.
+The four BugBench programs were selected as the only ones we could retrieve, build, and reproduce on our system.
 The BugsC++ suite does not include the original crash-causing inputs.
@@
 We randomly selected and reproduced four bugs from the “memory error” category.
-Some of the studied programs do not crash at run time.
+Some of the programs studied do not crash at run time.
 We employed AddressSanitizer (Serebryany et al., 2012) to force a crash at the moment a memory violation occurs to trigger those defects.
@@
 RQ3: Is ChatDBG effective at diagnosing and fixing bugs in unmanaged code (C/C++)?
-We ran our C/C++ experiments on an an x86 Ubuntu 22.04 server.
+We ran our C/C++ experiments on an x86 Ubuntu 22.04 server.
 We used Clang and LLDB 17 to compile and debug, using flags -g -Og -fno-omit-frame-pointer.
@@
 Figure 8.
-ChatDBG success rate at fixing the proximate or root cause in C/C++ programs.
+ChatDBG success rate at fixing the proximate or root cause in C/C++ programs (§5.2).
 ChatDBG successfully identified and fixed the root cause 36% of the time and the proximate cause an additional 55% of the time.
@@
 LLMs are also inherently stochastic, and it is possible to obtain unusually good results by chance.
-To mitigate this threat, our evaluation runs ChatDBG on each test program at least ten times.
+To mitigate this threat, our evaluation runs ChatDBG on each test program at least ten times, which produced stable and repeatable results with only small variation in aggregate.
 Our evaluation depends on a manual evaluation of whether ChatDBG’s explanation of a bug and its proposed fix are satisfactory.
-We mitigated the risks of subjectivity by using precisely-defined criteria decided upon in advance.
+We mitigated the risks of subjectivity by using precisely defined criteria decided upon in advance.
 Python fixes were deemed successful if the resulting code met the correctness requirements outlined in the assignment.
@@
 Our evaluation shows that engaging in a debugging dialog with ChatDBG can significantly assist in identifying root causes of errors and developing correct fixes.
+Data-Availability Statement
+ChatDBG is available on GitHub at github.com/plasma-umass/ChatDBG.

### v5

@@
 For the Python programs, a single query led to an actionable bug fix 67% of the time; one additional follow-up query increased the success rate to 85%.
-ChatDBG has seen rapid uptake; it has already been downloaded more than 65,000 times.
+ChatDBG has seen rapid uptake; it has already been downloaded more than 75,000 times.
 Debugging, Artificial Intelligence, Software Engineering
-$c$s: Computing methodologies Artificial intelligenceccs: Software and its engineering Software testing and debugging
+$c$s: Computing methodologies Artificial intelligenceccs: Software and its engineering Software testing and debuggingcopyright: ccdoi: 10.1145/3729355journalyear: 2025journal: PACMSEjournalvolume: 2journalnumber: FSEarticle: FSE085publicationmonth: 7
 1. Introduction
@@
 data breakpoints); inspecting local variables, globals, heap objects, and backtraces of the call stack; and resuming program execution line-by-line (single-step) or at the granularity of function calls.
-Debuggers can be helpful, but finding and fixing software defects remains a deeply challenging and time-consuming task (Zeller, 2009a; Eisenstadt, 1993; Layman et al., 2013).
+Debuggers can be helpful, but finding and fixing software defects remains a deeply challenging and time-consuming task (Zeller, 2009; Eisenstadt, 1993; Layman et al., 2013).
 Programmers must still reason about program behavior to ascertain what went wrong.
@@
 Specifically, the LLM issues “function calls” (OpenAI, 2023a) to run commands in the underlying debugger to investigate program state, execute code, or obtain source code.
-The results of those calls are sent back to the LLM for use while constructing its response.
+The results of those calls are sent back to the LLM to use in constructing its response.
 After answering a query, control is returned to the programmer, who may then enter additional commands or chat messages.
@@
 Our evaluation presents a range of case studies demonstrating that ChatDBG improves significantly on existing debuggers.
-On a suite of unpublished Python scripts and Jupyter notebooks written by undergraduate students, one or two queries is sufficient for ChatDBG to properly diagnose and fix defects 87% of the time, typically at a cost well under \$0.20 USD.
+On a suite of unpublished Python scripts and Jupyter notebooks written by undergraduate students, one or two queries is sufficient for ChatDBG to properly diagnose and fix defects 85% of the time, typically at a cost well under \$0.20 USD.
 ChatDBG is also effective at identifying causes and providing fixes for a range of real-world bugs in C/C++ code.
-This paper makes the following contributions.
+This paper makes the following contributions:
   • It introduces ChatDBG, an AI-powered debugger assistant that enables large language models to “take the wheel” and control the debugger via agentic reasoning.
@@
 ChatDBG is also useful for experienced programmers, who can augment debugging sessions with ChatDBG’s reasoning capabilities in a conversational and interactive way.
-This paper is organized as follows.
-Section 2 illustrates the use of ChatDBG to debug a program.
-Section 3 describes key related work.
-Section 4 describes ChatDBG’s implementation.
-Section 5 presents our evaluation.
-Section 6 outlines directions for future work, and Section 7 concludes.
 Source code for bootstrap.py
@@
 It is supposed to create an array of marble colors, compute the proportions of blue marbles in resamples of that array, and assert that their mean is about 0.7, the proportion for the array.
+[…]
+A Python code snippet.
 \$ chatdbg bootstrap.py Traceback (most recent call last):   File "/python3.11/site-packages/ipdb/__main__.py", ...
@@
 Commands issued by ChatDBG when it takes the wheel are highlighted in yellow, and its response is highlighted in blue.
+[…]
+Text snippet of a ChatDBG debugging session.
 2. Overview
@@
 ChatDBG identifies high variance resulting from the small number of trials as the root cause.
-ChatDBG identifies the root cause: proportion_blue fails to compute and return the desired statistic, and it provides a corrected version of proportion_blue.
+[…]
+Text snippet of a ChatDBG debugging session.
+ChatDBG identifies and corrects the root cause: proportion_blue incorrectly computes the desired statistic.
 When ChatDBG cannot identify the root cause, it suggests further debugging steps and control is returned to the user, who may continue the chat, issue further debugger commands, or both.
 Figure 3 illustrates this scenario, where a version of bootstrap.py with the corrected proportion_blue function fails the assertion on line 21.
-The user asks why the mean of stats is not close to 0.7, and ChatDBG’s first response suggests examining whether 0.7 is the appropriate expected value.
-The user then computes the proportion of blue marbles with a debugger command and tells ChatDBG that 0.7 is the correct value.
-In response, ChatDBG points to the low number of trials (5) as the issue.
+The user asks why the mean of stats is not close to 0.7, and ChatDBG’s initial response suggests examining whether 0.7 is the appropriate expected value.
+To verify this, the user then computes the proportion of blue marbles with a debugger command and tells ChatDBG that 0.7 is indeed the correct value.
+In its follow-up, ChatDBG points to the low number of trials (five) as the issue.
 The LLM drew this correct conclusion without seeing any discussion of trial size or variance in any program state, code, or documentation encountered during the chat.
@@
 3. Related Work
+Table 1.
+Debugger features and their dates of introduction (§3).
+Most key features have been around for decades.
+By integrating into modern debuggers (GDB, LLDB, and Pdb), ChatDBG inherits all of their features while significantly extending them with functionality to explain bugs and their root causes, propose fixes, and answer arbitrary natural-language queries over program state.
+(An asterisk or year in italics means the feature is limited in functionality, performance, or depends on specific hardware support.)
 System and Date Single Step Stack Navigation Breakpoints (BPs) Conditional BPs Source Level Trace Display State Eval. Code Watchpoints Explain Bugs Propose Fixes
@@
 ChatDBG, 2023
-Table 1.
-Debugger features and their dates of introduction (§3).
-Most key features have been around for decades.
-By integrating into modern debuggers (GDB, LLDB, and Pdb), ChatDBG inherits all of their features while significantly extending them with functionality to explain bugs and their root causes, propose fixes, and answer arbitrary natural-language queries over program state.
-(An asterisk or year in italics means the feature is limited in functionality, performance, or depends on specific hardware support.)
 Table 1 presents an overview of previous interactive debuggers, together with their features.
@@
 Figure 4. ChatDBG architecture and command processing algorithm (§4.3).
+[…]
+A graph showing the architecture of ChatDBG, along with a short algorithm summary.
+Users may directly issue standard debugger commands, that are handled by the existing debugger.
+They may also use free-form text, that will be interpreted by ChatDBG.
+ChatDBG will act as an agent, and facilitate conversation between the LLM and the underlying debugger.
+The LLM will issue commands to retrieve context about source code or runtime state before issuing its final response and/or recommendation.
 4.3. ChatDBG Architecture Overview
@@
 Why doesn’t stats have 5 elements?
+[…]
 Figure 5. The initial prompt for the debugging session in Figure 24.4). For brevity, the enriched stack includes only five lines of source in each frame, rather than the default of 10.
@@
 However, that approach seems less useful in this context, as program state information will be distinct for each debugging session and not easily vectorized.
+Table 2.
+ChatDBG command extensions (§4.6).
+ChatDBG extends the underlying debuggers with several new commands to help the LLM navigate through and understand the target’s code.
+ChatDBG provides access to them via the LLM’s function call API.
 Command Debugger Output
@@
 definition loc symbol LLDB The declaration for the first occurrence of symbol at loc, where loc has the form filename:lineno.
-Table 2.
-ChatDBG command extensions (§4.6).
-ChatDBG extends the underlying debuggers with several new commands to help the LLM navigate through and understand the target’s code.
-ChatDBG provides access to them via the LLM’s function call API.
 4.6. Navigating the Code
@@
 ChatDBG augments Pdb with the info command, which prints the docstring for any function, class, field, method, or package.
-It additionally prints the source for any user-defined code.
+It additionally prints the source code for any user-defined function.
 The info requests in Figure 2 demonstrate these two cases for proportion_blue and bootstrap_statistic, respectively.
-The command is implemented via the standard inspect and pydoc libraries.
+The command is implemented via the standard inspect and pydoc Python libraries.
 The info command is not directly reproducible for unmanaged code in LLVM because there is no comparable existing debugger support for retrieving the source or documentation for a symbol.
@@
 It returns the code for cells in that slice.
-Suppose the code from bootstrap.py in Figure 1 were written in four notebook cells:
+Suppose the code from bootstrap.py in Figure 1 were written in four notebook cells as shown below:
 […]
@@
 Instead, ChatDBG presents them to the user, who may then vet and judge the quality of the LLM’s responses and decide whether or not to follow suggested changes.
+Table 3.
+Python programs exhibiting a variety of common errors (§5.1).
+Programs c1–c8 are command line scripts, and programs s1–s14 are Jupyter notebooks, which utilize two non-standard libraries consisting of 3,000 lines of code.
+Semantic errors reflect failed tests expressed as assertions.
+Crashes reflect unexpected termination due to any other type of error.
 Name LoC Type Reported Exception Root Cause
@@
 c8 72 semantic Assertion Error Missing test for lowercase words
-s1 123 semantic Assertion Error Incorrect drop and rename operations leading to bad data in table
+s1 123 semantic Assertion Error Incorrect drop and rename operations leading to bad data
 s2 124 semantic Assertion Error Incorrect max operation on a table
@@
 s10 118 semantic Assertion Error Loops doesn’t append to array correctly
-s11 181 crash Value Error Creates a sample without replacement that is larger than the input
+s11 181 crash Value Error Creates a sample without replacement larger than the input
 s12 127 crash Value Error Incorrect label when accessing column value for table row
 s13 127 crash Value Error Pivot uses wrong columns for row/columns in new table
-s14 65 crash Index Error Incorrect computation of a random sample under null hypothesis
-Table 3.
-Python programs exhibiting a variety of common errors (§5.1).
-Programs c1–c8 are command line scripts, and programs s1–s14 are Jupyter notebooks, which utilize two non-standard libraries consisting of 3,000 lines of code.
-Semantic errors reflect failed tests expressed as assertions.
-Crashes reflect unexpected termination due to any other type of error.
+s14 65 crash Index Error Incorrect computation of random sample under null hypothesis
 5. Evaluation
@@
 5.1. Python
-We applied ChatDBG to all bugs in a collection of student labs from two introductory computer science courses; see Table 3.
+We applied ChatDBG to all of the bugs in a collection of student labs from two introductory computer science courses; see Table 3.
 Bugs c1–c8 are in non-interactive scripts from a programming class that perform various file reading and text processing tasks.
@@
 ChatDBG used the gpt-4-1106-preview model for these experiments.
-Under +Targeted Question, the first prompt and response led to, on average, a chat of about 10,000 tokens (7,500 words), a cost of about \$$0.120.12$ USD under OpenAI’s current pricing model (OpenAI, [n.
-d.]), and a completion time of about 25 seconds.
+Under +Targeted Question, the first prompt and response led to, on average, a chat of about 10,000 tokens (7,500 words), a cost of about \$$0.120.12$ USD under OpenAI’s current pricing model (OpenAI, 2024), and a completion time of about 25 seconds.
 Subsequent steps in extended debugging dialogs incurred comparable costs.
@@
 These characteristics will be different for other platforms and models and, given current trends, we expect significant reductions in both time and cost as models improve.
+Table 4. Configurations used in the Python experiments (§5.1).
 Configuration Stack Take the Initial Ask a
@@
 +Dialog enriched specialized
-Table 4. Configurations used in the Python experiments (§5.1).
 Figure 6. Overall ChatDBG success rate for each configuration (§5.1). ChatDBG innovations and user-provided context gradually increase effectiveness.
+[…]
+Chart showing the overall success rate of ChatDBG under each configuration. Each feature addition improves the success rate, culminating at 85% using the fully-featured ChatDBG.
 RQ1: Is ChatDBG effective at diagnosing and fixing bugs in Python?
@@
 Figure 7. Success rate for ChatDBG for each program and configuration (§5.1). Vertical lines show the mean.
-Error Fix
-Program LoC Type Root Cause Proximate Cause Root Cause
-BC (Lu et al., 2005) 17.0k Buffer overflow Input from data file printed to a fixed-size buffer Truncate on copy Use dynamic size
-GZIP (Lu et al., 2005) 8.2k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
-NCOM (Lu et al., 2005) 1.9k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
-PEG (An et al., 2023) 14.7k Null dereference Invalid input produces corrupted data structure Check if not null Warn/exit
-POLY (Lu et al., 2005) 0.7k Buffer overflow Command line argument is unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
-TIFF (An et al., 2023) 58.9k Division by zero Combination of command line options leads to a division by zero Override option values Warn/exit when invalid
-YAML1 (An et al., 2023) 8.7k Stack overflow Long sequences of { in the input leads to deep recursion Use iterative method Guard recursion depth
-YAML2 (An et al., 2023) 8.7k Assertion failure Specific input causes a peek request for non-existent “next” token Replace assert Check before peeking
-Table 5. Bugs in unmanaged C/C++ code, and our criteria for fixing the proximate cause or the root cause of each (§5.2).
+[…]
+A detailed version of Figure 6, showing the success rate for each program evaluated under each configuration. While there is some variation, the ChatDBG generally improves with additional features.
 The +Targeted Question configuration demonstrates the impact of providing even the most modest details about expected behavior in queries.
@@
 We attribute this result to the fact that longer and more complex prompts can occasionally degrade the effectiveness of LLMs (Levy et al., 2024).
-In general, the further the distance between the root cause of a bug and observable effect, the more challenging it was for ChatDBG (and people (Zeller, 2009b, p.
+In general, the further the distance between the root cause of a bug and observable effect, the more challenging it was for ChatDBG (and people (Zeller, 2009, p.
 243)) to find it.
@@
 We randomly selected and reproduced four bugs from the “memory error” category.
+Table 5. Bugs in unmanaged C/C++ code, and our criteria for fixing the proximate cause or the root cause of each (§5.2).
+Error Fix
+Program LoC Type Root Cause Proximate Cause Root Cause
+BC (Lu et al., 2005) 17.0k Buffer overflow Input from data file printed to a fixed-size buffer Truncate on copy Use dynamic size
+GZIP (Lu et al., 2005) 8.2k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
+NCOM (Lu et al., 2005) 1.9k Buffer overflow Command line argument unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
+PEG (An et al., 2023) 14.7k Null dereference Invalid input produces corrupted data structure Check if not null Warn/exit
+POLY (Lu et al., 2005) 0.7k Buffer overflow Command line argument is unsafely copied to a fixed-size buffer Truncate on copy Check size & warn/exit
+TIFF (An et al., 2023) 58.9k Division by zero Combination of command line options leads to a division by zero Override option values Warn/exit when invalid
+YAML1 (An et al., 2023) 8.7k Stack overflow Long sequences of { in the input leads to deep recursion Use iterative method Guard recursion depth
+YAML2 (An et al., 2023) 8.7k Assertion failure Specific input causes a peek request for non-existent “next” token Replace assert Check before peeking
 Some of the programs studied do not crash at run time.
@@
 While fixing root causes is the ultimate goal, fixing proximate causes can still be beneficial as fixing crashes enables further debugging steps.
-Figure 8.
-ChatDBG success rate at fixing the proximate or root cause in C/C++ programs (§5.2).
-ChatDBG successfully identified and fixed the root cause 36% of the time and the proximate cause an additional 55% of the time.
 Figure 8 presents ChatDBG’s ability to suggest a fix for either the proximate or root cause of the bug.
@@
 It was also capable of providing relevant, actionable fixes: 36% of its suggestions addressed the root cause of the bug, while another 55% corrected the proximate cause.
+Figure 8.
+ChatDBG success rate at fixing the proximate or root cause in C/C++ programs (§5.2).
+ChatDBG successfully identified and fixed the root cause 36% of the time and the proximate cause an additional 55% of the time.
+[…]
+Chart showing the success rate of ChatDBG at fixing the proximate or root cause in C/C++ programs.
+Proximate causes are identified and fixed most of the time, while success finding and fixing root cause varies among programs.
 5.3. Threats to Validity
@@
 Our evaluation shows that engaging in a debugging dialog with ChatDBG can significantly assist in identifying root causes of errors and developing correct fixes.
+Acknowledgements.
+This material is based upon work supported by the National Science Foundation under Grant No.
+2243636.
+Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
 Data-Availability Statement
-ChatDBG is available on GitHub at github.com/plasma-umass/ChatDBG.
+ChatDBG is available on GitHub at github.com/plasma-umass/ChatDBG. An archived version is also available on Zenodo (Levin et al., 2025).